### v1

# ms-Mamba: Multi-scale Mamba for Time-Series Forecasting

Abstract
The problem of Time-series Forecasting is generally addressed by recurrent, Transformer-based and the recently proposed Mamba-based architectures. However, existing architectures generally process their input at a single temporal scale, which may be sub-optimal for many tasks where information changes over multiple time scales. In this paper, we introduce a novel architecture called Multi-scale Mamba (ms-Mamba) to address this gap. ms-Mamba incorporates multiple temporal scales by using multiple Mamba blocks with different sampling rates ($\Delta$s). Our experiments on many benchmarks demonstrate that ms-Mamba outperforms state-of-the-art approaches, including the recently proposed Transformer-based and Mamba-based models. Codes and models will be made available.
[…]
XXX
1 Introduction
Time-series Forecasting (TSF) is the problem of predicting future values of a variable of interest, given its history. This fundamental problem used to be generally addressed using recurrent architectures [37, 11] and long-short term memory [19] or their variants [7, 14], see, e.g., [24, 27] for detailed surveys. Such models are inherently well-suited to the task due to their sequential information modeling abilities. The introduction of self-attention based architectures, a.k.a. Transformers [32], enabled attending to more informative patterns and correlations across time and provided significant improvements. However, Transformers’ quadratic computational complexity has been a limiting factor.
State Space Models (SSMs) [17, 31] are reported to provide a better balance between performance and computational complexity. Recently proposed architectures based on SSMs, namely, Mamba [16], offer the promise of on-par or better performance than Transformer-based alternatives while running significantly faster. This has led to the widespread use of Mamba or its derivatives across different domains [16, 43, 30, 41].
The use of a Mamba-based approach for TSF was recently explored by Wang et al.[34]. Wang et al.introduced an architecture, called S-Mamba, which used Mamba in both the forward and reverse directions for TSF. Wang et al.showed that this simple approach obtained state-of-the-art (SOTA) results on many TSF benchmarks, often providing significant gains over Transformer-based architectures.
Figure 1: (a) Mamba and its variations (S-Mamba) use a single time-scale while processing time-series data. (b) Our ms-Mamba processes its input at different time-scales to better capture signal at different scales.
Time-series data generally consist of signals of multiple temporal scales. To better capture and exploit the multi-scale nature of time-series data, the literature has introduced extensions over the conventional models; e.g., multi-scale recurrent architectures [8], multi-scale convolution [22] and multi-scale Transformers [47, 5].
In this paper, we introduce ms-Mamba, a Mamba-based architecture for multi-scale processing of time-series data. To be specific, by leveraging on the versatility of SSMs’ learnable sampling rate, we construct a block that consists of multiple SSMs with different independent or inter-related sampling rates. We show that our ms-Mamba performs better than Transformer-based and Mamba-based architectures on several datasets.
Contributions. Our main contributions are as follows:
  • We propose a multi-scale architecture based on Mamba. To do so, we use multiple SSMs with different sampling rates to process the signal at different temporal scales.
  • We introduce and compare different strategies for using different sampling rates for different SSMs: (1) Using hyper-parameters as multipliers for a learned sampling rate, (2) learning different sampling rate for each SSM, (3) and estimating sampling rates from the input.
  • We show that, on the commonly used TSF benchmarks, our ms-Mamba surpasses or performs on par with SOTA models. For example, on the Solar-Energy dataset, ms-Mamba outperforms its closest competitor s-Mamba (0.229 vs. 0.240 in terms of mean-squared error) with less parameters, less memory footprint and less computational overhead.
2 Related Work
2.1 Time-series Forecasting
2.1.1 Transformer-based Models
Transformers, initially introduced by Vaswani et al.[32], have revolutionized tasks that involve sequence processing and generation, with their self-attention mechanism proving highly effective in capturing long-range dependencies. This architecture, originally designed for natural language processing, has since been adapted to time-series forecasting tasks [1], primarily due to its ability to model complex temporal relationships. Duong-Trung et al.[10] demonstrate the efficacy of Transformers in long-term multi-horizon forecasting, addressing the challenge of vanishing correlations over extended horizons.
Recent studies have aimed to address the limitations of standard Transformers in time-series applications. Foumani et al.[12] propose enhanced positional encodings to improve the positional awareness of the Transformer architecture in multivariate time-series classification. Lim et al.[25] propose a Transformer architecture to make use of a complex mix of inputs. Wang et al.[33] present Graphformer, a model that replaces traditional convolutional layers with dilated convolutional layers, thereby improving the efficiency of capturing temporal dependencies across multiple variates in a graph-based framework.
Despite their unprecedented successes in natural language processing tasks, Transformer models face several challenges when applied to other time-series domains. One key limitation is their content-based attention mechanism, which struggles to detect crucial temporal dependencies, particularly in cases where dependencies weaken over time or when strong seasonal patterns are present [38]. Additionally, Transformers suffer from the quadratic complexity of the attention mechanism, which increases computational costs and memory usage, for long input sequences [36].
To address these issues, several studies have proposed modifications to the self-attention mechanism. For instance, Zhou et al.[48] introduce Informer that employs a sparsified self-attention operation to lower computational complexity and improve long-term forecasting efficiency. Similarly, Wu et al.[39] propose Autoformer, which relies on an auto-correlation-based self-attention to better capture temporal dependencies.
2.1.2 Linear Models
Linear models are another popular approach in TSF due to their simplicity and efficiency [2]. Oreshkin et al.[29] propose a stacked MLP based architecture with residual links. Challu et al.[3] improve this architecture with multi-rate data sampling and hierarchical interpolation for effectively modeling extra long sequences. Zeng et al.[44] analyze Transformers for TSF and found that simple linear mappings can outperform Transformer models especially when the data has strong periodic patterns. Chen et al.[6] introduce another notable linear approach, TSMixer, which leverages an all-MLP architecture to efficiently incorporate cross-variate and auxiliary information. Zhang et al.[45] propose LightTS which is tailored towards efficiently handling very long input series in multivariate TSF. Wang et al.[35] propose time-series Multi-layer Perceptron (MLP), which improves forecasting performance by incorporating domain-specific knowledge into the MLP architecture.
While linear models with MLPs are simpler architectures and faster compared to Transformer-based models, they face several limitations. These models generally struggle with non-linear dependencies and tend to underperform in scenarios involving highly volatile or non-stationary patterns [6]. Moreover, compared to Transformer-based models, linear architectures are less effective at capturing global dependencies. This limitation necessitates longer input sequences to achieve comparable forecasting performance, which can increase the computational cost [42].
2.2 Mamba Models
Mamba [16] is a recent sequence model based on State Space Models (SSMs) [17, 31]. Due to its promise of better efficiency-performance trade-off, Mamba has quickly attracted interest from researchers across different domains [30, 41]. Mamba’s ability to perform content-based reasoning in linear complexity to the sequence length with its hardware-aware algorithm, made it an attractive alternative to the Transformer models. Several works have explored its application to time-series forecasting tasks. Wang et al.[34] propose S-Mamba, which relies on a bidirectional Mamba layer to capture inter-variate dependencies and an MLP to extract temporal dependencies. Their model achieves SOTA performance while being faster than Transformer-based alternatives.
Figure 2: An overview of the proposed method. ms-Mamba processes the time-series data at different sampling rates to better capture the multi-scale nature of the input signal. This is achieved by processing and updating the embeddings with different sampling rates (SR).
3 Preliminaries and Background
3.1 Problem Formulation
Time-series forecasting is the problem of estimating future $F$ values $\mathbf{Y}_{t+1:t+F}=\{\mathbf{x}_{t+1},\mathbf{x}_{t+2},...,\mathbf{x}_{t+F}\}\in\mathbb{R}^{F\times D}$ of a multi-variate time-series data given its recent $L$ values $\mathbf{X}_{t-L:t}=\{\mathbf{x}_{t-L},\mathbf{x}_{t-L+1},...,\mathbf{x}_{t}\}\in\mathbb{R}^{L\times D}$ as input. The task is to find the mapping $f$:
$f:\mathbf{X}_{t-L:t}\rightarrow\mathbf{Y}_{t+1:t+F},$ (1)
which is represented by a deep network whose parameters are estimated from a training dataset.
3.2 State Space Models (SSMs)
SSMs [17, 31] are sequence models which use a latent state space representation for representing a mapping between a time-series input $x(t)$ and the output $y(t)$ (considering a single-variate setting to simplify notation):
$\delta h(t)$ $=$ $\textbf{A}h(t)+\textbf{B}x(t),$ (2)
$y(t)$ $=$ $\textbf{C}h(t),$ (3)
where $\mathbf{A},\mathbf{B},\mathbf{C}$ are matrices with learnable values; $h(t)$ is the latent (state) representation; and $\delta h(t)$ is the update for the latent space with the current input, $x(t)$. This continuous formulation is transformed into a discrete model with a sampling rate $\Delta$ as follows:
$h_{t}$ $=$ $\hat{\textbf{A}}h_{t-1}+\hat{\textbf{B}}x_{t},$ (4)
$y_{t}$ $=$ ${\textbf{C}}h_{t},$ (5)
where $\hat{\textbf{A}},\hat{\textbf{B}},{\textbf{C}}$ are derived by the chosen sampling function (e.g., for zeroth-order hold sampling, $\hat{\textbf{A}}=\exp(\Delta\mathbf{A})$, $\hat{\mathbf{B}}=(\Delta\mathbf{A})^{-1}(\exp(\Delta\mathbf{A}-I)\cdot\Delta\mathbf{B}$ [16]).
SSMs have been recently extended to work more efficiently through the use of a convolutional approach [18] and more effectively through better initialization [13]. Moreover, by relating $\hat{\textbf{A}},\hat{\textbf{B}},{\textbf{C}}$ to the input, the Mamba model [16] provides comparable or better performance than its Transformer-based counterparts.
4 Methodology: ms-Mamba
In this section, we describe our ms-Mamba in detail. For multi-scale temporal processing, ms-Mamba essentially leverages multiple Mamba blocks with different $\Delta$ working in parallel. See Figure 2 for an overview.
4.1 Embedding Layer
Following prior work [15, 26], we first transform the input time-series data through an embedding layer (Figure 2). Given an input sequence $\mathbf{X}\in\mathbb{R}^{L\times D}$ with $L$ time steps and $D$ variables, we apply a linear transformation along the temporal dimension:
$\mathbf{E}=\mathrm{Embedding}(\mathbf{X})\in\mathbb{R}^{D_{e}\times D},$ (6)
where $D_{e}$ is the embedding dimension. This transformation maps each time-series from length $L$ to length $D_{e}$ while preserving the number of variables $D$, thus enabling us to deal with fixed-length tokens instead of variable input sequence length $L$.
4.2 Multi-scale Mamba Layer
As summarized in Section 3.2, SSMs, Mamba and their variants process time at one learnable sampling rate, $\Delta$. Our architecture ms-Mamba addresses this gap by processing the input at different sampling rates $\Delta_{1}$, $\Delta_{2}$, …, $\Delta_{s}$. This is achieved by combining multiple Mamba blocks with different sampling rates as follows:
$\mathbf{E}^{l}_{m}=\mathrm{Avg}(\mathrm{Mamba}(\mathbf{E}^{l};\Delta_{1}),...,\mathrm{Mamba}(\mathbf{E}^{l};\Delta_{n})),$ (7)
where $\mathbf{E}^{l}$ is the output of the embedding layer at layer $l$. We explore three different strategies for obtaining $\Delta_{i}$:
  1. 1.
    Fixed temporal scales, where $\Delta_{1}$ is kept learnable (as in the original Mamba model) but $\Delta_{2},\Delta_{3},...,\Delta_{n}$ are taken as multiples of $\Delta_{1}$:
    $\Delta_{i}=\alpha_{i}\times\Delta_{1},\quad\quad i\in\{2,...,n\},$ (8)
    where $\alpha_{i}$ are hyper-parameters.
  2. 2.
    Learnable temporal scales, where all $\Delta_{i}$ are defined as learnable variables as in the original Mamba model.
  3. 3.
    Dynamic temporal scales, where all $\Delta_{i}$ are estimated through a Multi-layer Perceptron:
    $\Delta_{i}=\mathrm{MLP}(\mathrm{Flatten}(\mathbf{E}^{l})),$ (9)
    where $\mathrm{Flatten}(\cdot)$ reshapes the input tensor $\mathbf{E}^{l}\in\mathbb{R}^{L\times D_{e}}$ into a vector of dimension $L\cdot D_{e}$, and $\mathrm{MLP}(\cdot)$ consists of two linear layers with a ReLU activation in between: $\mathrm{MLP}(\mathbf{x})=\mathbf{W}_{2}\max(0,\mathbf{W}_{1}\mathbf{x}+\mathbf{b}_{1})+\mathbf{b}_{2}$, mapping the flattened input to $n$ different sampling rates.
To improve the effectiveness of sequential processing, we employ our Multi-scale Mamba module in both directions as illustrated in Figure 2, following prior work (e.g., [15, 50]).
4.3 Normalization, Feed-Forward Network and Projection
The output of the Multi-scale Mamba Layer ($\mathbf{E}^{l}_{m}$) passes through Layer Normalization, a multi-layer perceptron (MLP – with two layers with the ReLU nonlinearity) to obtain the embeddings for the next layer ($\mathbf{E}^{l+1}\in\mathbb{R}^{L\times D_{e}}$):
$\mathbf{E}^{l+1}=\mathrm{MLP}(\mathrm{LayerNorm}(\mathbf{E}^{l}_{m})).$ (10)
After the last encoder block ($\mathbf{E}^{L}\in\mathbb{R}^{L\times D_{e}}$), a linear projection layer is applied to map the embedding dimension to prediction length to obtain the final prediction ($\hat{y}\in\mathbb{R}^{F\times D}$):
$\hat{y}=\mathrm{Linear(\mathbf{E}^{L})}.$ (11)
4.4 Training Objective
The model is trained to minimize the Mean Square Error (MSE) between the predicted values and the ground truth:
$\mathcal{L}=\frac{1}{F\times D}\sum_{i=1}^{F}\sum_{j=1}^{D}(\hat{y}_{i,j}-y_{i,j})^{2},$ (12)
where $\hat{y}_{i,j}$ and $y_{i,j}$ are the predicted and ground truth values for the $i$-th time step and $j$-th variable, respectively; $F$ is the forecast horizon; and $D$ is the number of variables. The model parameters are optimized using the Adam optimizer [20] – see the Suppl. Mat. for more details about the training and experimental details.
5 Experiments
5.1 Experimental Details
Datasets. To evaluate our proposed ms-Mamba, we conduct extensive experiments on thirteen real-world time-series forecasting benchmark datasets. The datasets are grouped into three categories for easier comparison. (1) Traffic-related datasets, which include Traffic [39] and PEMS [4]. The Traffic dataset consists of hourly road occupancy rates from the California Department of Transportation, consisting of data collected from 862 sensors on San Francisco Bay area freeways from January 2015 to December 2016. PEMS datasets are complex spatial-temporal datasets for California’s public traffic networks, includes four subsets (PEMS03, PEMS04, PEMS07, PEMS08), similar to SCINet. These traffic-related datasets has many periodic features. (2) ETT (Electricity Transformer Temperature) datasets [48], which contain load and oil temperature data from electricity transformers, collected between July 2016 and July 2018. This group includes four subsets: ETTm1, ETTm2, ETTh1, and ETTh2, which have fewer variables and show less regularity compared to traffic datasets. (3) Other datasets: Electricity [39], Exchange [39], Weather [39], and Solar-Energy [21]. The Electricity dataset includes the hourly electricity usage of 321 customers from 2012 to 2014. Solar-Energy dataset contains solar power generation data from 137 PV plants in Alabama in 2006, recorded at 10 minute resolution. The Weather dataset includes 21 meteorological indicators also recorded at 10 minute resolution from the Max Planck Biogeochemistry Institute’s Weather Station in 2020. Exchange dataset compiles daily exchange rates for eight countries from 1990 to 2016. The prior two datasets of this category contain many features most of which are periodic, the last two datasets have fewer primarily aperiodic features. See the Suppl. Mat for a summary of the datasets.
Models ms-Mamba ms-Mamba w/ Learnable $\Delta$ S-Mamba iTrans- former RLinear PatchTST Cross- former TiDE TimesNet DLinear FED- former Auto- former
Traffic 96 0.376 0.375 0.382 0.395 0.649 0.462 0.522 0.805 0.593 0.650 0.587 0.613
192 0.392 0.384 0.396 0.417 0.601 0.466 0.530 0.756 0.617 0.598 0.604 0.616
336 0.405 0.408 0.417 0.433 0.609 0.482 0.558 0.762 0.629 0.605 0.621 0.622
720 0.452 0.442 0.460 0.467 0.647 0.514 0.589 0.719 0.640 0.645 0.626 0.660
Avg 0.406 0.402 0.414 0.428 0.626 0.481 0.550 0.760 0.620 0.625 0.610 0.628
PEMS03 12 0.065 0.066 0.065 0.071 0.126 0.099 0.090 0.178 0.085 0.122 0.126 0.272
24 0.087 0.087 0.087 0.093 0.246 0.142 0.121 0.257 0.118 0.201 0.149 0.334
48 0.131 0.133 0.133 0.125 0.551 0.211 0.202 0.379 0.155 0.333 0.227 1.032
96 0.197 0.201 0.201 0.164 1.057 0.269 0.262 0.490 0.228 0.457 0.348 1.031
Avg 0.120 0.122 0.122 0.113 0.495 0.180 0.169 0.326 0.147 0.278 0.213 0.667
PEMS04 12 0.074 0.072 0.076 0.078 0.138 0.105 0.098 0.219 0.087 0.148 0.138 0.424
24 0.086 0.083 0.084 0.095 0.258 0.153 0.131 0.292 0.10 0.224 0.177 0.459
48 0.102 0.099 0.115 0.120 0.572 0.229 0.205 0.409 0.136 0.355 0.270 0.646
96 0.130 0.121 0.137 0.150 1.137 0.291 0.402 0.492 0.190 0.452 0.341 0.912
Avg 0.098 0.094 0.103 0.111 0.526 0.195 0.209 0.353 0.129 0.295 0.231 0.610
PEMS07 12 0.060 0.060 0.063 0.067 0.118 0.095 0.094 0.173 0.082 0.115 0.109 0.199
24 0.075 0.075 0.081 0.088 0.242 0.150 0.139 0.271 0.101 0.210 0.125 0.323
48 0.091 0.091 0.093 0.110 0.562 0.253 0.311 0.446 0.134 0.398 0.165 0.390
96 0.111 0.109 0.117 0.139 1.096 0.346 0.396 0.628 0.181 0.594 0.262 0.554
Avg 0.085 0.084 0.089 0.101 0.504 0.211 0.235 0.380 0.124 0.329 0.165 0.367
PEMS08 12 0.074 0.073 0.076 0.079 0.133 0.168 0.165 0.227 0.112 0.154 0.173 0.436
24 0.097 0.098 0.104 0.115 0.249 0.224 0.215 0.318 0.141 0.248 0.210 0.467
48 0.156 0.154 0.167 0.186 0.569 0.321 0.315 0.497 0.198 0.440 0.320 0.966
96 0.243 0.236 0.245 0.221 1.166 0.408 0.377 0.721 0.320 0.674 0.442 1.385
Avg 0.143 0.140 0.148 0.150 0.529 0.280 0.268 0.441 0.193 0.379 0.286 0.814
Table 1: Experiment 1: Quantitative comparison between ms-Mamba and the existing methods on traffic-related datasets. The lookback length $L$ is set to 96 and the forecast length $T$ is set to 12, 24, 48, 96 for PEMS and 96, 192, 336, 720 for Traffic. Top results are highlighted in bold while the second bests are underlined.
Compared Methods. We compare our model with 10 state-of-the-art (SOTA) time-series forecasting models belonging to 4 different model families: (1) Mamba based models: S-Mamba [34]; (2) Transformer based models: iTransformer [26], PatchTST [28], Crossformer [46], FEDformer [49], Autoformer [39]; (3) Linear based models: TiDE [9], DLinear [44], RLinear [23]; and (4) Temporal Convolution based models: TimesNet [40]. The following provides a brief overview of these models:
  • S-Mamba [34] employs a bidirectional Mamba encoder block to capture inter-variate correlations and a feed forward network temporal dependency encoding layer to learn temporal sequence dynamics. This novel approach is the current SOTA model for TSF task and forms the foundation of our proposed method.
  • iTransformer [26] inverts the order of sequence processing by first analyzing each individual variate separately and then merging the information across all variates.
  • PatchTST [28] divides the time-series into sub-series patches treated as input tokens. It leverages channel-independent shared embeddings and weights for efficient representation learning.
  • Crossformer [46] embeds multivariate time-series into a 2D vector array preserving time and dimension information and introduces two-stage attention to capture both cross-time and cross-dimension dependencies.
  • FEDformer [49] is a frequency-enhanced Transformer that uses trend-seasonality decomposition and exploit sparse representations, Fourier transform, of time-series data to achieve linear complexity to sequence length.
  • Autoformer [39] constructs a decomposition architecture that employs traditional sequence decomposition in its inner blocks and utilizes an auto-correlation mechanism.
  • DLinear [44] maps trend and seasonality components into predictions via a single linear layer.
  • TiDE [9] is a MLP based encoder-decoder model that is best suitable for linear dynamical systems.
  • RLinear [23] is the current SOTA linear model that introduces reversible normalization and channel independence within a purely linear structure.
  • TimesNet [40] proposes a task-general backbone, TimesBlock, that transforms 1D time-series into 2D tensors and uses 2D convolution kernels to capture intra-period and inter-period variations.
Training and Implementation Details. See the Suppl Mat for training and implementation details, especially the hyperparameters.
Performance Measure. Following the common practice (e.g., [34, 26, 28]), models’ performances are compared using the Mean Square Error (MSE) as defined in Equation (12).
Models ms-Mamba ms-Mamba w/ Learnable $\Delta$ S-Mamba iTrans- former RLinear PatchTST Cross- former TiDE TimesNet DLinear FED- former Auto- former
ETTm1 96 0.328 0.326 0.333 0.334 0.355 0.329 0.404 0.364 0.338 0.345 0.379 0.505
192 0.372 0.372 0.376 0.377 0.391 0.367 0.450 0.398 0.374 0.380 0.426 0.553
336 0.406 0.406 0.408 0.426 0.424 0.399 0.532 0.428 0.410 0.413 0.445 0.621
720 0.470 0.470 0.475 0.491 0.487 0.454 0.666 0.487 0.478 0.474 0.543 0.671
Avg 0.394 0.394 0.398 0.407 0.414 0.387 0.513 0.419 0.400 0.403 0.448 0.588
ETTm2 96 0.176 0.175 0.179 0.180 0.182 0.175 0.287 0.207 0.187 0.193 0.203 0.255
192 0.244 0.244 0.250 0.250 0.246 0.241 0.414 0.290 0.249 0.284 0.269 0.281
336 0.309 0.306 0.312 0.311 0.307 0.305 0.597 0.377 0.321 0.369 0.325 0.339
720 0.408 0.407 0.411 0.412 0.407 0.402 1.730 0.558 0.408 0.554 0.421 0.433
Avg 0.284 0.283 0.288 0.288 0.286 0.281 0.757 0.358 0.291 0.350 0.305 0.327
ETTh1 96 0.384 0.384 0.386 0.386 0.386 0.414 0.423 0.479 0.384 0.386 0.376 0.449
192 0.437 0.438 0.443 0.441 0.437 0.460 0.471 0.525 0.435 0.436 0.420 0.500
336 0.479 0.482 0.489 0.487 0.479 0.501 0.570 0.565 0.491 0.481 0.459 0.521
720 0.482 0.493 0.502 0.503 0.481 0.500 0.653 0.594 0.521 0.519 0.506 0.514
Avg 0.445 0.449 0.455 0.454 0.446 0.469 0.529 0.541 0.458 0.456 0.440 0.496
ETTh2 96 0.291 0.291 0.296 0.297 0.288 0.302 0.745 0.400 0.340 0.333 0.358 0.346
192 0.371 0.369 0.376 0.380 0.374 0.388 0.877 0.528 0.402 0.477 0.429 0.456
336 0.411 0.412 0.424 0.428 0.415 0.426 1.043 0.643 0.452 0.594 0.496 0.482
720 0.418 0.418 0.426 0.427 0.420 0.431 1.104 0.874 0.462 0.831 0.463 0.515
Avg 0.373 0.373 0.381 0.383 0.374 0.387 0.942 0.611 0.414 0.559 0.437 0.450
Table 2: Experiment 1: Quantitative comparison between ms-Mamba and the existing methods on ETT Datasets. The lookback length $L$ is set to 96 and the forecast length $T$ is set to 96, 192, 336, 720. Top results are highlighted in bold while the second bests are underlined.
Models ms-Mamba ms-Mamba w/ Learnable $\Delta$ S-Mamba iTrans- former RLinear PatchTST Cross- former TiDE TimesNet DLinear FED- former Auto- former
Electricity 96 0.137 0.138 0.139 0.148 0.201 0.181 0.219 0.237 0.168 0.197 0.193 0.201
192 0.157 0.157 0.159 0.162 0.201 0.188 0.231 0.236 0.184 0.196 0.201 0.222
336 0.171 0.174 0.176 0.178 0.215 0.204 0.246 0.249 0.198 0.209 0.214 0.231
720 0.195 0.199 0.204 0.225 0.257 0.246 0.280 0.284 0.220 0.245 0.246 0.254
Avg 0.165 0.167 0.170 0.178 0.219 0.205 0.244 0.251 0.192 0.212 0.214 0.227
Exchange 96 0.085 0.086 0.086 0.086 0.093 0.088 0.256 0.094 0.107 0.088 0.148 0.197
192 0.177 0.178 0.182 0.177 0.184 0.176 0.470 0.184 0.226 0.176 0.271 0.300
336 0.325 0.326 0.332 0.331 0.351 0.301 1.268 0.349 0.367 0.313 0.460 0.509
720 0.843 0.843 0.867 0.847 0.886 0.901 1.767 0.852 0.964 0.839 1.195 1.447
Avg 0.358 0.358 0.367 0.360 0.378 0.367 0.940 0.370 0.416 0.354 0.519 0.613
Weather 96 0.163 0.163 0.165 0.174 0.192 0.177 0.158 0.202 0.172 0.196 0.217 0.266
192 0.213 0.213 0.214 0.221 0.240 0.225 0.206 0.242 0.219 0.237 0.276 0.307
336 0.269 0.270 0.274 0.278 0.292 0.278 0.272 0.287 0.280 0.283 0.339 0.359
720 0.349 0.349 0.350 0.358 0.364 0.354 0.398 0.351 0.365 0.345 0.403 0.419
Avg 0.249 0.249 0.251 0.258 0.272 0.259 0.259 0.271 0.259 0.265 0.309 0.338
Solar Ener. 96 0.196 0.195 0.205 0.203 0.322 0.234 0.310 0.312 0.250 0.290 0.242 0.884
192 0.230 0.230 0.237 0.233 0.359 0.267 0.734 0.339 0.296 0.320 0.285 0.834
336 0.250 0.247 0.258 0.248 0.397 0.290 0.750 0.368 0.319 0.353 0.282 0.941
720 0.249 0.249 0.260 0.249 0.397 0.289 0.769 0.370 0.338 0.356 0.357 0.882
Avg 0.231 0.229 0.240 0.233 0.369 0.270 0.641 0.347 0.301 0.330 0.291 0.885
Table 3: Experiment 1: Quantitative comparison between ms-Mamba and the existing methods on Electricity, Exchange, Weather and Solar-Energy Datasets. The lookback length $L$ is set to 96 and the forecast length $T$ is set to 96, 192, 336, 720. Top results are highlighted in bold while the second bests are underlined.
5.2 Experiment 1: Comparison with State-of-the-Art
We compare ms-Mamba with fixed and learnable temporal scales against SOTA methods over 13 benchmark datasets.
Traffic-related Datasets (Table 1). The results in Table 1 show that ms-Mamba with fixed or learnable temporal scales provides the best or second best results over all traffic datasets across all forecast lengths. Compared to our baseline model of S-Mamba, ms-Mamba delivers significant improvements, especially on the Traffic dataset.
ETT Datasets (Table 2). On ETT datasets, as in Table 2, ms-Mamba with fixed or learnable temporal scales is typically the second best and the best performing method in ETTh2 and in some configurations of ETTm1 and ETTm2. Compared to S-Mamba, ms-Mamba provides better performance.
Other Datasets (Table 3). The results in other datasets (Table 3) are in agreement with the results on Traffic-related and ETT datasets: our ms-Mamba with fixed or learnable temporal scales performs better than or is generally on par with SOTA methods, and consistently provides better performance than the S-Mamba baseline.
                                    Dataset $\Rightarrow$ Traffic ETTh1 Solar Energy
                             VC Encoding $\Rightarrow$ 96 192 336 720 96 192 336 720 96 192 336 720
ms-Mamba with fixed scales: $\mathbf{\alpha}=(1,2,4,8)$ 0.376 0.392 0.405 0.452 0.384 0.437 0.479 0.482 0.196 0.230 0.250 0.249
ms-Mamba with fixed scales: $\mathbf{\alpha}=(0.5,1,1.5,2)$ 0.374 0.389 0.414 0.458 0.386 0.438 0.481 0.491 0.197 0.232 0.248 0.251
ms-Mamba with fixed scales: $\mathbf{\alpha}=(1,2,3,4)$ 0.390 0.403 0.415 0.455 0.386 0.439 0.482 0.495 0.196 0.232 0.250 0.251
ms-Mamba with fixed scales: $\mathbf{\alpha}=(1,4,8,16)$ 0.380 0.411 0.421 0.453 0.385 0.439 0.484 0.495 0.197 0.232 0.250 0.251
ms-Mamba with learnable scales 0.375 0.384 0.408 0.442 0.384 0.438 0.482 0.493 0.195 0.230 0.247 0.249
ms-Mamba with dynamic scales 0.376 0.390 0.414 0.454 0.384 0.440 0.480 0.493 0.194 0.231 0.249 0.251
Table 4: Experiment 2: Ablation study on Traffic, ETTh1 and Solar Energy datasets (one dataset from each category). The lookback length $L=96$, while the forecast length $T\in\left\{96,192,336,720\right\}$. $\mathbf{\alpha}=(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4})$ indicates that the $\Delta_{1}$ (learnable sampling rate of the base Mamba) is multiplied by these coefficients to obtain the sampling rates for the Mamba blocks. Top results are highlighted in bold while the second bests are underlined.
5.3 Experiment 2: Ablation Analysis
In this experiment, we investigate the different strategies for integrating multiple-scales into ms-Mamba. For this, we evaluate the performances of the strategies described in Section 4.2: (i) ms-Mamba with fixed temporal scales, where we multiply the learnable sampling rate $\Delta_{1}$ by fixed hyperparameters, $\Delta_{i}=\alpha_{i}\times\Delta_{1}$. (ii) ms-Mamba with learnable temporal scales, where each $\Delta_{i}$ is a learnable parameter. (iii) ms-Mamba with dynamic temporal scales, where each $\Delta_{i}$ is estimated by an MLP applied on the input embeddings. For this analysis, we tune the hyperparameters in all settings. To keep the number of experiments manageable, we consider one dataset from each category.
The results in Table 4 suggest that using fixed temporal scales with $\alpha_{i}$ coefficients of $(1,2,4,8)$ performs best among different coefficients considered (6 best and 2 second-best results out of 12 datasets). Smaller or larger coefficients than $(1,2,4,8)$ do not appear to provide better results overall. Moreover, ms-Mamba with learnable $\Delta_{i}$ provides slightly better results than the best fixed temporal scale version (provides 6 best and 4 second-best results).
Fixed temporal scales introduces several hyperparameters to be tuned, which is a significant limitation. Therefore, ms-Mamba with learnable temporal scales can be the preferred option. However, in Experiment 1, we have considered both versions for a more extensive evaluation.
Models ms-Mamba S-Mamba
Metric MSE #Params Memory MACs MSE #Params Memory MACs
ETTh2 96 0.291 0.481M 1.84MB 0.165G 0.296 1.150M 4.40MB 1.563G
192 0.369 0.484M 1.85MB 0.171G 0.376 1.175M 4.48MB 1.580G
336 0.412 0.503M 1.92MB 0.180G 0.424 1.212M 4.62MB 1.606G
720 0.418 0.552M 2.11MB 0.195G 0.426 1.311M 5.00MB 1.675G
Solar Ener. 96 0.195 3.958M 15.10MB 16.72G 0.205 4.643M 17.71MB 20.00G
192 0.230 2.028M 7.74MB 8.57G 0.237 4.692M 17.90MB 20.21G
336 0.247 4.015M 15.31MB 16.99G 0.258 4.766M 18.18MB 20.54G
720 0.249 4.113M 15.70MB 17.43G 0.260 4.963M 18.93MB 21.40G
Traffic 96 0.375 29.68M 113.2MB 403.5G 0.382 9.186M 35.04MB 125.0G
192 0.384 14.94M 56.99MB 203.1G 0.396 9.236M 35.23MB 125.7G
336 0.408 29.81M 113.7MB 405.2G 0.417 9.310M 35.51MB 126.7G
720 0.442 29.68M 114.5MB 407.9G 0.460 9.507M 36.26MB 129.5G
Table 5: Experiment 3: Performance comparison of ms-Mamba and S-Mamba on one dataset from each category. The lookback length $L$ is set to 96 and the forecast length $T$ is set to 96, 192, 336, 720.
5.4 Experiment 3: Efficiency Analysis
In this experiment, we investigate the efficiency of ms-Mamba (with learnable temporal scales, as it provides the best results and incurs more learnable parameters) in comparison with the baseline S-Mamba. In each dataset and for each method, we provide the results of the best performing configurations.
As listed in Table 5, on the ETTh1 and Solar Energy datasets, we see that ms-Mamba provides the best results with less parameters, memory and operations. This is crucial as it shows multiple temporal scales can be utilized with less computational overhead. However, this result is not observed in the Traffic dataset because it contains significantly more variates (862) compared to the ETTh2 (7) and Solar Energy (137) datasets. Although ms-Mamba provides better results with more variates, it is not able to do so with less parameters, memory and computations as in ETTh2 and Solar Energy datasets.
6 Conclusion
In this paper, we introduce a novel multi-scale architecture for the problem of time-series forecasting (TSF). Our architecture extends Mamba (or its derivative S-Mamba) where we include several Mamba blocks with different sampling rates to process multiple temporal scales simultaneously. The different sampling rates can either be fixed or learned from the data, which leads to a simple architecture with a multi-scale ability. Our results on several TSF benchmarks show that our approach provides the best or on par performance compared to SOTA methods. What is remarkable is that, compared to the baseline model (S-Mamba), ms-Mamba provides better results with less parameters, memory, and operations.
Limitations and Future Work. It is a promising research direction to apply ms-Mamba for other types of modalities, e.g., text and images. Moreover, ms-Mamba can complement other types of deep modules, e.g., scaled-dot-product attention, linear attention.

### v2

@@
-ms-Mamba: Multi-scale Mamba for Time-Series Forecasting
+ms-Mamba: Multi-scale Mamba for Time-Series Forecasting11footnotemark: 1
 Abstract
-The problem of Time-series Forecasting is generally addressed by recurrent, Transformer-based and the recently proposed Mamba-based architectures.
-However, existing architectures generally process their input at a single temporal scale, which may be sub-optimal for many tasks where information changes over multiple time scales.
-In this paper, we introduce a novel architecture called Multi-scale Mamba (ms-Mamba) to address this gap.
-ms-Mamba incorporates multiple temporal scales by using multiple Mamba blocks with different sampling rates ($\Delta$s).
-Our experiments on many benchmarks demonstrate that ms-Mamba outperforms state-of-the-art approaches, including the recently proposed Transformer-based and Mamba-based models.
-Codes and models will be made available.
 […]
-XXX
+keywords:
+Time-series forecasting , Mamba , Multi-scale Mamba
+journal: Neurocomputing
+[…]
+[ceng]organization=Dept. of Computer Eng. and ROMER Robotics Center, Middle East Technical University, state=Ankara, country=Turkey
+[…]
+[arch]organization=Dept. of Architecture and ROMER Robotics Center, Middle East Technical University, state=Ankara, country=Turkey
 1 Introduction
-Time-series Forecasting (TSF) is the problem of predicting future values of a variable of interest, given its history.
-This fundamental problem used to be generally addressed using recurrent architectures [37, 11] and long-short term memory [19] or their variants [7, 14], see, e.g., [24, 27] for detailed surveys.
+Time-series Forecasting (TSF) is the problem of predicting future values of a variable of interest, given its history (Lim and Zohren, 2021; Miller et al., 2024; Nobrega and Oliveira, 2019; George et al., 2023).
+This fundamental problem used to be generally addressed using recurrent architectures (Williams and Zipser, 1989; Elman, 1990) and long-short term memory (Hochreiter and Schmidhuber, 1997) or their variants (Chung et al., 2014; Graves and Schmidhuber, 2005), see, e.g., (Lim and Zohren, 2021; Miller et al., 2024) for detailed surveys.
 Such models are inherently well-suited to the task due to their sequential information modeling abilities.
 The introduction of self-attention based architectures, a.k.a.
-Transformers [32], enabled attending to more informative patterns and correlations across time and provided significant improvements.
+Transformers (Vaswani et al., 2017), enabled attending to more informative patterns and correlations across time and provided significant improvements.
 However, Transformers’ quadratic computational complexity has been a limiting factor.
-State Space Models (SSMs) [17, 31] are reported to provide a better balance between performance and computational complexity.
-Recently proposed architectures based on SSMs, namely, Mamba [16], offer the promise of on-par or better performance than Transformer-based alternatives while running significantly faster.
-This has led to the widespread use of Mamba or its derivatives across different domains [16, 43, 30, 41].
-The use of a Mamba-based approach for TSF was recently explored by Wang et al.[34].
-Wang et al.introduced an architecture, called S-Mamba, which used Mamba in both the forward and reverse directions for TSF.
-Wang et al.showed that this simple approach obtained state-of-the-art (SOTA) results on many TSF benchmarks, often providing significant gains over Transformer-based architectures.
-Figure 1: (a) Mamba and its variations (S-Mamba) use a single time-scale while processing time-series data.
-(b) Our ms-Mamba processes its input at different time-scales to better capture signal at different scales.
-Time-series data generally consist of signals of multiple temporal scales.
-To better capture and exploit the multi-scale nature of time-series data, the literature has introduced extensions over the conventional models; e.g., multi-scale recurrent architectures [8], multi-scale convolution [22] and multi-scale Transformers [47, 5].
-In this paper, we introduce ms-Mamba, a Mamba-based architecture for multi-scale processing of time-series data.
+State Space Models (SSMs) (Gu et al., 2021a; Smith et al., 2022) are reported to provide a better balance between performance and computational complexity.
+Recently proposed architectures based on SSMs, namely, Mamba (Gu and Dao, 2023), offer the promise of on-par or better performance than Transformer-based alternatives while running significantly faster.
+This has led to the wide-spread use of Mamba or its derivatives across different domains (Gu and Dao, 2023; Yue and Li, 2024; Qu et al., 2024; Xu et al., 2024).
+The use of a Mamba-based approach for TSF was recently explored by Wang et al.
+(2024c).
+Wang et al.
+introduced an architecture, called S-Mamba, which used Mamba in both the forward and reverse directions for TSF.
+Wang et al.
+showed that this simple approach obtained state-of-the-art (SOTA) results on many TSF benchmarks, often providing significant gains over Transformer-based architectures.
+Figure 1: (a) Time-series data often contain information at multiple time-scales.
+Illustrated here is temperature from the Weather dataset Wu et al.
+(2021), averaged over different window lengths to highlight different scales.
+See B.3 for similar multi-scale visualizations from other datasets.
+(b) Mamba and its variations (S-Mamba) use a single time-scale while processing time-series data.
+(c) Our ms-Mamba processes its input at different time-scales to better capture signal at different scales.
+Motivation: Why multi-scale TSF?
+Time-series data generally consist of signals of multiple temporal scales (see, e.g., Figure 1(a)).
+A temperature signal, for example, can have trends at scales of hours, days (day and night), weeks, months or years.
+To better capture and exploit the multi-scale nature of time-series data, the literature has introduced extensions over the conventional models; e.g., multi-scale recurrent architectures (Chung et al., 2017), multi-scale convolution (Li et al., 2024) and multi-scale Transformers (Zhang et al., 2024; Chen et al., 2024).
+Our Solution.
+In this paper, we introduce ms-Mamba, a Mamba-based architecture for multi-scale processing of time-series data (Figure 1(c)), in contrast to standard Mamba-based models that process data at a single time-scale (Figure 1(b)).
 To be specific, by leveraging on the versatility of SSMs’ learnable sampling rate, we construct a block that consists of multiple SSMs with different independent or inter-related sampling rates.
@@
 Contributions. Our main contributions are as follows:
+  • 1.
     We propose a multi-scale architecture based on Mamba. To do so, we use multiple SSMs with different sampling rates to process the signal at different temporal scales.
+  • 2.
     We introduce and compare different strategies for using different sampling rates for different SSMs: (1) Using hyper-parameters as multipliers for a learned sampling rate, (2) learning different sampling rate for each SSM, (3) and estimating sampling rates from the input.
+  • 3.
     We show that, on the commonly used TSF benchmarks, our ms-Mamba surpasses or performs on par with SOTA models.
@@
 2.1 Time-series Forecasting
-2.1.1 Transformer-based Models
-Transformers, initially introduced by Vaswani et al.[32], have revolutionized tasks that involve sequence processing and generation, with their self-attention mechanism proving highly effective in capturing long-range dependencies.
-This architecture, originally designed for natural language processing, has since been adapted to time-series forecasting tasks [1], primarily due to its ability to model complex temporal relationships.
-Duong-Trung et al.[10] demonstrate the efficacy of Transformers in long-term multi-horizon forecasting, addressing the challenge of vanishing correlations over extended horizons.
+Transformer-based Models.
+Transformers, initially introduced by Vaswani et al.
+(2017), have revolutionized tasks that involve sequence processing and generation, with their self-attention mechanism proving highly effective in capturing long-range dependencies.
+This architecture, originally designed for natural language processing, has since been adapted to time-series forecasting tasks (Ahmed et al., 2023), primarily due to its ability to model complex temporal relationships.
+Duong-Trung et al.
+(2023) demonstrate the efficacy of Transformers in long-term multi-horizon forecasting, addressing the challenge of vanishing correlations over extended horizons.
 Recent studies have aimed to address the limitations of standard Transformers in time-series applications.
-Foumani et al.[12] propose enhanced positional encodings to improve the positional awareness of the Transformer architecture in multivariate time-series classification.
-Lim et al.[25] propose a Transformer architecture to make use of a complex mix of inputs.
-Wang et al.[33] present Graphformer, a model that replaces traditional convolutional layers with dilated convolutional layers, thereby improving the efficiency of capturing temporal dependencies across multiple variates in a graph-based framework.
+Foumani et al.
+(2024) propose enhanced positional encodings to improve the positional awareness of the Transformer architecture in multivariate time-series classification.
+Lim et al.
+(2021) propose a Transformer architecture to make use of a complex mix of inputs.
+Wang et al.
+(2024a) present Graphformer, a model that replaces traditional convolutional layers with dilated convolutional layers, thereby improving the efficiency of capturing temporal dependencies across multiple variates in a graph-based framework.
 Despite their unprecedented successes in natural language processing tasks, Transformer models face several challenges when applied to other time-series domains.
-One key limitation is their content-based attention mechanism, which struggles to detect crucial temporal dependencies, particularly in cases where dependencies weaken over time or when strong seasonal patterns are present [38].
-Additionally, Transformers suffer from the quadratic complexity of the attention mechanism, which increases computational costs and memory usage, for long input sequences [36].
+One key limitation is their content-based attention mechanism, which struggles to detect crucial temporal dependencies, particularly in cases where dependencies weaken over time or when strong seasonal patterns are present (Woo et al., 2022).
+Additionally, Transformers suffer from the quadratic complexity of the attention mechanism, which increases computational costs and memory usage, for long input sequences (Wen et al., 2022).
 To address these issues, several studies have proposed modifications to the self-attention mechanism.
-For instance, Zhou et al.[48] introduce Informer that employs a sparsified self-attention operation to lower computational complexity and improve long-term forecasting efficiency.
-Similarly, Wu et al.[39] propose Autoformer, which relies on an auto-correlation-based self-attention to better capture temporal dependencies.
-2.1.2 Linear Models
-Linear models are another popular approach in TSF due to their simplicity and efficiency [2].
-Oreshkin et al.[29] propose a stacked MLP based architecture with residual links.
-Challu et al.[3] improve this architecture with multi-rate data sampling and hierarchical interpolation for effectively modeling extra long sequences.
-Zeng et al.[44] analyze Transformers for TSF and found that simple linear mappings can outperform Transformer models especially when the data has strong periodic patterns.
-Chen et al.[6] introduce another notable linear approach, TSMixer, which leverages an all-MLP architecture to efficiently incorporate cross-variate and auxiliary information.
-Zhang et al.[45] propose LightTS which is tailored towards efficiently handling very long input series in multivariate TSF.
-Wang et al.[35] propose time-series Multi-layer Perceptron (MLP), which improves forecasting performance by incorporating domain-specific knowledge into the MLP architecture.
+For instance, Zhou et al.
+(2021) introduce Informer that employs a sparsified self-attention operation to lower computational complexity and improve long-term forecasting efficiency.
+Similarly, Wu et al.
+(2021) propose Autoformer, which relies on an auto-correlation-based self-attention to better capture temporal dependencies.
+Linear Models.
+Linear models are another popular approach in TSF due to their simplicity and efficiency (Benidis et al., 2022).
+Oreshkin et al.
+(2019) propose a stacked MLP based architecture with residual links.
+Challu et al.
+(2023) improve this architecture with multi-rate data sampling and hierarchical interpolation for effectively modeling extra long sequences.
+Zeng et al.
+(2023) analyze Transformers for TSF and found that simple linear mappings can outperform Transformer models especially when the data has strong periodic patterns.
+Chen et al.
+(2023) introduce another notable linear approach, TSMixer, which leverages an all-MLP architecture to efficiently incorporate cross-variate and auxiliary information.
+Zhang et al.
+(2022) propose LightTS which is tailored towards efficiently handling very long input series in multivariate TSF.
+Wang et al.
+(2024b) propose time-series Multi-layer Perceptron (MLP), which improves forecasting performance by incorporating domain-specific knowledge into the MLP architecture.
 While linear models with MLPs are simpler architectures and faster compared to Transformer-based models, they face several limitations.
-These models generally struggle with non-linear dependencies and tend to underperform in scenarios involving highly volatile or non-stationary patterns [6].
+These models generally struggle with non-linear dependencies and tend to underperform in scenarios involving highly volatile or non-stationary patterns (Chen et al., 2023).
 Moreover, compared to Transformer-based models, linear architectures are less effective at capturing global dependencies.
-This limitation necessitates longer input sequences to achieve comparable forecasting performance, which can increase the computational cost [42].
+This limitation necessitates longer input sequences to achieve comparable forecasting performance, which can increase the computational cost (Yi et al., 2024).
+Multi-scale Models.
+The literature has witnessed many multi-scale time-series models using different architectures.
+For instance, exploiting the Transformer architecture, Pyraformer (Liu et al., 2022b) constructs a multi-resolution pyramidal graph to capture long-range dependencies, while Crossformer (Zhang and Yan, 2023) utilizes Dimension-Segment-Wise embeddings to explicitly model information at varying granularities.
+Beyond pure attention, TimesNet (Wu et al., ) transforms 1D series into 2D to capture intra- and inter-period variations, and convolutional approaches like MICN (Wang et al., 2023) and SCINet (Liu et al., 2022a) leverage multi-branch convolutions or recursive downsampling to fuse local and global contexts.
+While these methods achieve multi-scale processing through explicit structural hierarchies or downsampling operations, our approach extends this capability to Mamba by leveraging the inherent discretization properties of the sampling rate $\Delta$, enabling multi-resolution feature extraction without architectural downsampling.
 2.2 Mamba Models
-Mamba [16] is a recent sequence model based on State Space Models (SSMs) [17, 31].
-Due to its promise of better efficiency-performance trade-off, Mamba has quickly attracted interest from researchers across different domains [30, 41].
+Mamba Gu and Dao (2023) is a recent sequence model based on State Space Models (SSMs) Gu et al.
+(2021a); Smith et al.
+(2022).
+Due to its promise of better efficiency-performance trade-off, Mamba has quickly attracted interest from researchers across different domains Qu et al.
+(2024); Xu et al.
+(2024).
 Mamba’s ability to perform content-based reasoning in linear complexity to the sequence length with its hardware-aware algorithm, made it an attractive alternative to the Transformer models.
 Several works have explored its application to time-series forecasting tasks.
-Wang et al.[34] propose S-Mamba, which relies on a bidirectional Mamba layer to capture inter-variate dependencies and an MLP to extract temporal dependencies.
+Wang et al.
+(2024c) propose S-Mamba, which relies on a bidirectional Mamba layer to capture inter-variate dependencies and an MLP to extract temporal dependencies.
 Their model achieves SOTA performance while being faster than Transformer-based alternatives.
@@
 This is achieved by processing and updating the embeddings with different sampling rates (SR).
+In the input visualization, each square represents a data point for a single variate at a specific time step, with the axes representing the Variates and Time dimensions, respectively.
 3 Preliminaries and Background
 3.1 Problem Formulation
-Time-series forecasting is the problem of estimating future $F$ values $\mathbf{Y}_{t+1:t+F}=\{\mathbf{x}_{t+1},\mathbf{x}_{t+2},...,\mathbf{x}_{t+F}\}\in\mathbb{R}^{F\times D}$ of a multi-variate time-series data given its recent $L$ values $\mathbf{X}_{t-L:t}=\{\mathbf{x}_{t-L},\mathbf{x}_{t-L+1},...,\mathbf{x}_{t}\}\in\mathbb{R}^{L\times D}$ as input.
+Time-series forecasting is the problem of estimating future $T$ values $\mathbf{Y}_{t+1:t+T}=\{\mathbf{x}_{t+1},\mathbf{x}_{t+2},...,\mathbf{x}_{t+T}\}\in\mathbb{R}^{F\times D}$ of a multi-variate time-series data given its recent $L$ values $\mathbf{X}_{t-L:t}=\{\mathbf{x}_{t-L},\mathbf{x}_{t-L+1},...,\mathbf{x}_{t}\}\in\mathbb{R}^{L\times D}$ as input.
 The task is to find the mapping $f$:
-$f:\mathbf{X}_{t-L:t}\rightarrow\mathbf{Y}_{t+1:t+F},$ (1)
+$f:\mathbf{X}_{t-L:t}\rightarrow\mathbf{Y}_{t+1:t+T},$ (1)
 which is represented by a deep network whose parameters are estimated from a training dataset.
 3.2 State Space Models (SSMs)
-SSMs [17, 31] are sequence models which use a latent state space representation for representing a mapping between a time-series input $x(t)$ and the output $y(t)$ (considering a single-variate setting to simplify notation):
+SSMs Gu et al.
+(2021a); Smith et al.
+(2022) are sequence models which use a latent state space representation for representing a mapping between a time-series input $x(t)$ and the output $y(t)$ (considering a single-variate setting to simplify notation):
 $\delta h(t)$ $=$ $\textbf{A}h(t)+\textbf{B}x(t),$ (2)
@@
 $y_{t}$ $=$ ${\textbf{C}}h_{t},$ (5)
-where $\hat{\textbf{A}},\hat{\textbf{B}},{\textbf{C}}$ are derived by the chosen sampling function (e.g., for zeroth-order hold sampling, $\hat{\textbf{A}}=\exp(\Delta\mathbf{A})$, $\hat{\mathbf{B}}=(\Delta\mathbf{A})^{-1}(\exp(\Delta\mathbf{A}-I)\cdot\Delta\mathbf{B}$ [16]).
-SSMs have been recently extended to work more efficiently through the use of a convolutional approach [18] and more effectively through better initialization [13].
-Moreover, by relating $\hat{\textbf{A}},\hat{\textbf{B}},{\textbf{C}}$ to the input, the Mamba model [16] provides comparable or better performance than its Transformer-based counterparts.
+where $\hat{\textbf{A}},\hat{\textbf{B}},{\textbf{C}}$ are derived by the chosen sampling function (e.g., for zeroth-order hold sampling, $\hat{\textbf{A}}=\exp(\Delta\mathbf{A})$, $\hat{\mathbf{B}}=(\Delta\mathbf{A})^{-1}(\exp(\Delta\mathbf{A}-I)\cdot\Delta\mathbf{B}$ Gu and Dao (2023)).
+SSMs have been recently extended to work more efficiently through the use of a convolutional approach Gu et al.
+(2021b) and more effectively through better initialization Fu et al.
+(2022).
+Moreover, by relating $\hat{\textbf{A}},\hat{\textbf{B}},{\textbf{C}}$ to the input, the Mamba model Gu and Dao (2023) provides comparable or better performance than its Transformer-based counterparts.
 4 Methodology: ms-Mamba
@@
 See Figure 2 for an overview.
+Influence of Sampling Rate on Temporal Resolution.
+By employing multiple Mamba blocks, each parameterized with a different sampling rate ($\Delta$), our model is designed to intrinsically capture features at multiple, distinct temporal resolutions (or time scales): Overall, large $\Delta$ induces shorter memory and lower temporal resolution whereas small $\Delta$ facilitates long-memory, high-resolution behavior.
+See Section A for a more in-depth argument.
 4.1 Embedding Layer
-Following prior work [15, 26], we first transform the input time-series data through an embedding layer (Figure 2).
+Following prior work Grazzi et al.
+(2024); Liu et al.
+(2023), we first transform the input time-series data through an embedding layer (Figure 2).
 Given an input sequence $\mathbf{X}\in\mathbb{R}^{L\times D}$ with $L$ time steps and $D$ variables, we apply a linear transformation along the temporal dimension:
@@
 $\mathbf{E}^{l}_{m}=\mathrm{Avg}(\mathrm{Mamba}(\mathbf{E}^{l};\Delta_{1}),...,\mathrm{Mamba}(\mathbf{E}^{l};\Delta_{n})),$ (7)
-where $\mathbf{E}^{l}$ is the output of the embedding layer at layer $l$. We explore three different strategies for obtaining $\Delta_{i}$:
+where $\mathbf{E}^{l}$ is the output of the embedding layer at layer $l$.
+Note that each mamba block contains an internal normalization layer, so our averaging operation fuses the already-normalized outputs from each scale.
+Furthermore, keep in mind that the sampling rate $\Delta$ is an internal SSM parameter and not an input downsampling operation.
+All parallel Mamba blocks receive the same input $\mathbf{E}^{l}$ and produce identically-sized outputs, allowing for direct fusion without any temporal alignment.
+We explore three different strategies for obtaining $\Delta_{i}$:
   1. 1.
@@
     $\Delta_{i}=\mathrm{MLP}(\mathrm{Flatten}(\mathbf{E}^{l})),$ (9)
-    where $\mathrm{Flatten}(\cdot)$ reshapes the input tensor $\mathbf{E}^{l}\in\mathbb{R}^{L\times D_{e}}$ into a vector of dimension $L\cdot D_{e}$, and $\mathrm{MLP}(\cdot)$ consists of two linear layers with a ReLU activation in between: $\mathrm{MLP}(\mathbf{x})=\mathbf{W}_{2}\max(0,\mathbf{W}_{1}\mathbf{x}+\mathbf{b}_{1})+\mathbf{b}_{2}$, mapping the flattened input to $n$ different sampling rates.
-To improve the effectiveness of sequential processing, we employ our Multi-scale Mamba module in both directions as illustrated in Figure 2, following prior work (e.g., [15, 50]).
+    where $\mathrm{Flatten}(\cdot)$ reshapes the input tensor $\mathbf{E}^{l}\in\mathbb{R}^{L\times D_{e}}$ into a vector of dimension $L\cdot D_{e}$, and $\mathrm{MLP}(\cdot)$ consists of two linear layers with a ReLU activation in between: $\mathrm{MLP}(\mathbf{x})=\mathbf{W}_{2}\max(0,\mathbf{W}_{1}\mathbf{x}+\mathbf{b}_{1})+\mathbf{b}_{2}$, mapping the flattened input ($\mathbf{x}$, representing $\mathrm{Flatten}(\mathbf{E}^{l})$ in Equation 9) to $n$ different sampling rates.
+To improve the effectiveness of sequential processing, we employ our Multi-scale Mamba module in both directions as illustrated in Figure 2, following prior work (e.g., Grazzi et al.
+(2024); Zhu et al.
+(2024)).
+By using different sampling rates $\Delta_{i}$, each Mamba block generates different discrete-time SSM parameters, allowing each block to specialize in capturing temporal dependencies at a distinct scale.
+Section A for a more in-depth argument.
 4.3 Normalization, Feed-Forward Network and Projection
@@
 $\mathbf{E}^{l+1}=\mathrm{MLP}(\mathrm{LayerNorm}(\mathbf{E}^{l}_{m})).$ (10)
-After the last encoder block ($\mathbf{E}^{L}\in\mathbb{R}^{L\times D_{e}}$), a linear projection layer is applied to map the embedding dimension to prediction length to obtain the final prediction ($\hat{y}\in\mathbb{R}^{F\times D}$):
-$\hat{y}=\mathrm{Linear(\mathbf{E}^{L})}.$ (11)
+After the last ($N^{th}$) encoder block ($\mathbf{E}^{N}\in\mathbb{R}^{L\times D_{e}}$), a linear projection layer is applied to map the embedding dimension to prediction length to obtain the final prediction ($\hat{y}\in\mathbb{R}^{F\times D}$):
+$\hat{y}=\mathrm{Linear(\mathbf{E}^{N})}.$ (11)
 4.4 Training Objective
@@
 where $\hat{y}_{i,j}$ and $y_{i,j}$ are the predicted and ground truth values for the $i$-th time step and $j$-th variable, respectively; $F$ is the forecast horizon; and $D$ is the number of variables.
-The model parameters are optimized using the Adam optimizer [20] – see the Suppl.
+The model parameters are optimized using the Adam optimizer Kingma (2014) – see the Suppl.
 Mat.
@@
 5 Experiments
+Unless otherwise stated, ms-Mamba refers to ms-Mamba with learnable temporal scales.
 5.1 Experimental Details
@@
 The datasets are grouped into three categories for easier comparison.
-(1) Traffic-related datasets, which include Traffic [39] and PEMS [4].
+(1) Traffic-related datasets, which include Traffic Wu et al.
+(2021) and PEMS Chen et al.
+(2001).
 The Traffic dataset consists of hourly road occupancy rates from the California Department of Transportation, consisting of data collected from 862 sensors on San Francisco Bay area freeways from January 2015 to December 2016.
 PEMS datasets are complex spatial-temporal datasets for California’s public traffic networks, includes four subsets (PEMS03, PEMS04, PEMS07, PEMS08), similar to SCINet.
-These traffic-related datasets has many periodic features.
-(2) ETT (Electricity Transformer Temperature) datasets [48], which contain load and oil temperature data from electricity transformers, collected between July 2016 and July 2018.
+These traffic-related datasets have many periodic features.
+(2) ETT (Electricity Transformer Temperature) datasets Zhou et al.
+(2021), which contain load and oil temperature data from electricity transformers, collected between July 2016 and July 2018.
 This group includes four subsets: ETTm1, ETTm2, ETTh1, and ETTh2, which have fewer variables and show less regularity compared to traffic datasets.
-(3) Other datasets: Electricity [39], Exchange [39], Weather [39], and Solar-Energy [21].
+(3) Other datasets: Electricity Wu et al.
+(2021), Exchange Wu et al.
+(2021), Weather Wu et al.
+(2021), and Solar-Energy Lai et al.
+(2018).
 The Electricity dataset includes the hourly electricity usage of 321 customers from 2012 to 2014.
@@
 The prior two datasets of this category contain many features most of which are periodic, the last two datasets have fewer primarily aperiodic features.
-See the Suppl.
-Mat for a summary of the datasets.
-Models ms-Mamba ms-Mamba w/ Learnable $\Delta$ S-Mamba iTrans- former RLinear PatchTST Cross- former TiDE TimesNet DLinear FED- former Auto- former
-Traffic 96 0.376 0.375 0.382 0.395 0.649 0.462 0.522 0.805 0.593 0.650 0.587 0.613
-192 0.392 0.384 0.396 0.417 0.601 0.466 0.530 0.756 0.617 0.598 0.604 0.616
-336 0.405 0.408 0.417 0.433 0.609 0.482 0.558 0.762 0.629 0.605 0.621 0.622
-720 0.452 0.442 0.460 0.467 0.647 0.514 0.589 0.719 0.640 0.645 0.626 0.660
-Avg 0.406 0.402 0.414 0.428 0.626 0.481 0.550 0.760 0.620 0.625 0.610 0.628
-PEMS03 12 0.065 0.066 0.065 0.071 0.126 0.099 0.090 0.178 0.085 0.122 0.126 0.272
-24 0.087 0.087 0.087 0.093 0.246 0.142 0.121 0.257 0.118 0.201 0.149 0.334
-48 0.131 0.133 0.133 0.125 0.551 0.211 0.202 0.379 0.155 0.333 0.227 1.032
-96 0.197 0.201 0.201 0.164 1.057 0.269 0.262 0.490 0.228 0.457 0.348 1.031
-Avg 0.120 0.122 0.122 0.113 0.495 0.180 0.169 0.326 0.147 0.278 0.213 0.667
-PEMS04 12 0.074 0.072 0.076 0.078 0.138 0.105 0.098 0.219 0.087 0.148 0.138 0.424
-24 0.086 0.083 0.084 0.095 0.258 0.153 0.131 0.292 0.10 0.224 0.177 0.459
-48 0.102 0.099 0.115 0.120 0.572 0.229 0.205 0.409 0.136 0.355 0.270 0.646
-96 0.130 0.121 0.137 0.150 1.137 0.291 0.402 0.492 0.190 0.452 0.341 0.912
-Avg 0.098 0.094 0.103 0.111 0.526 0.195 0.209 0.353 0.129 0.295 0.231 0.610
-PEMS07 12 0.060 0.060 0.063 0.067 0.118 0.095 0.094 0.173 0.082 0.115 0.109 0.199
-24 0.075 0.075 0.081 0.088 0.242 0.150 0.139 0.271 0.101 0.210 0.125 0.323
-48 0.091 0.091 0.093 0.110 0.562 0.253 0.311 0.446 0.134 0.398 0.165 0.390
-96 0.111 0.109 0.117 0.139 1.096 0.346 0.396 0.628 0.181 0.594 0.262 0.554
-Avg 0.085 0.084 0.089 0.101 0.504 0.211 0.235 0.380 0.124 0.329 0.165 0.367
-PEMS08 12 0.074 0.073 0.076 0.079 0.133 0.168 0.165 0.227 0.112 0.154 0.173 0.436
-24 0.097 0.098 0.104 0.115 0.249 0.224 0.215 0.318 0.141 0.248 0.210 0.467
-48 0.156 0.154 0.167 0.186 0.569 0.321 0.315 0.497 0.198 0.440 0.320 0.966
-96 0.243 0.236 0.245 0.221 1.166 0.408 0.377 0.721 0.320 0.674 0.442 1.385
-Avg 0.143 0.140 0.148 0.150 0.529 0.280 0.268 0.441 0.193 0.379 0.286 0.814
-Table 1: Experiment 1: Quantitative comparison between ms-Mamba and the existing methods on traffic-related datasets.
+See the Suppl. Mat for more details on the datasets.
+Table 1: Experiment 1: Quantitative results on traffic-related datasets.
 The lookback length $L$ is set to 96 and the forecast length $T$ is set to 12, 24, 48, 96 for PEMS and 96, 192, 336, 720 for Traffic.
 Top results are highlighted in bold while the second bests are underlined.
+Dataset $T$ ms-Mamba (Ours) S-Mamba iTransformer RLinear PatchTST Crossformer TiDE TimesNet DLinear FEDformer Autoformer
+Traffic 96 0.375 0.382 0.395 0.649 0.462 0.522 0.805 0.593 0.650 0.587 0.613
+192 0.384 0.396 0.417 0.601 0.466 0.530 0.756 0.617 0.598 0.604 0.616
+336 0.408 0.417 0.433 0.609 0.482 0.558 0.762 0.629 0.605 0.621 0.622
+720 0.442 0.460 0.467 0.647 0.514 0.589 0.719 0.640 0.645 0.626 0.660
+[.4pt/1pt] Avg 0.402 0.414 0.428 0.626 0.481 0.550 0.760 0.620 0.625 0.610 0.628
+PEMS03 12 0.066 0.065 0.071 0.126 0.099 0.090 0.178 0.085 0.122 0.126 0.272
+24 0.087 0.087 0.093 0.246 0.142 0.121 0.257 0.118 0.201 0.149 0.334
+48 0.133 0.133 0.125 0.551 0.211 0.202 0.379 0.155 0.333 0.227 1.032
+96 0.201 0.201 0.164 1.057 0.269 0.262 0.490 0.228 0.457 0.348 1.031
+[.4pt/1pt] Avg 0.122 0.122 0.113 0.495 0.180 0.169 0.326 0.147 0.278 0.213 0.667
+PEMS04 12 0.072 0.076 0.078 0.138 0.105 0.098 0.219 0.087 0.148 0.138 0.424
+24 0.083 0.084 0.095 0.258 0.153 0.131 0.292 0.10 0.224 0.177 0.459
+48 0.099 0.115 0.120 0.572 0.229 0.205 0.409 0.136 0.355 0.270 0.646
+96 0.121 0.137 0.150 1.137 0.291 0.402 0.492 0.190 0.452 0.341 0.912
+[.4pt/1pt] Avg 0.094 0.103 0.111 0.526 0.195 0.209 0.353 0.129 0.295 0.231 0.610
+PEMS07 12 0.060 0.063 0.067 0.118 0.095 0.094 0.173 0.082 0.115 0.109 0.199
+24 0.075 0.081 0.088 0.242 0.150 0.139 0.271 0.101 0.210 0.125 0.323
+48 0.091 0.093 0.110 0.562 0.253 0.311 0.446 0.134 0.398 0.165 0.390
+96 0.109 0.117 0.139 1.096 0.346 0.396 0.628 0.181 0.594 0.262 0.554
+[.4pt/1pt] Avg 0.084 0.089 0.101 0.504 0.211 0.235 0.380 0.124 0.329 0.165 0.367
+PEMS08 12 0.073 0.076 0.079 0.133 0.168 0.165 0.227 0.112 0.154 0.173 0.436
+24 0.098 0.104 0.115 0.249 0.224 0.215 0.318 0.141 0.248 0.210 0.467
+48 0.154 0.167 0.186 0.569 0.321 0.315 0.497 0.198 0.440 0.320 0.966
+96 0.236 0.245 0.221 1.166 0.408 0.377 0.721 0.320 0.674 0.442 1.385
+[.4pt/1pt] Avg 0.140 0.148 0.150 0.529 0.280 0.268 0.441 0.193 0.379 0.286 0.814
 Compared Methods.
-We compare our model with 10 state-of-the-art (SOTA) time-series forecasting models belonging to 4 different model families: (1) Mamba based models: S-Mamba [34]; (2) Transformer based models: iTransformer [26], PatchTST [28], Crossformer [46], FEDformer [49], Autoformer [39]; (3) Linear based models: TiDE [9], DLinear [44], RLinear [23]; and (4) Temporal Convolution based models: TimesNet [40].
+We compare our model with 10 state-of-the-art (SOTA) time-series forecasting models belonging to 4 different model families: (1) Mamba based models: S-Mamba Wang et al.
+(2024c); (2) Transformer based models: iTransformer Liu et al.
+(2023), PatchTST Nie et al.
+(2022), Crossformer Zhang and Yan (2023), FEDformer Zhou et al.
+(2022), Autoformer Wu et al.
+(2021); (3) Linear based models: TiDE Das et al.
+(2023), DLinear Zeng et al.
+(2023), RLinear Li et al.
+(2023); and (4) Temporal Convolution based models: TimesNet Wu et al.
+.
 The following provides a brief overview of these models:
-  • S-Mamba [34] employs a bidirectional Mamba encoder block to capture inter-variate correlations and a feed forward network temporal dependency encoding layer to learn temporal sequence dynamics.
+  • 1.
+    S-Mamba Wang et al.
+(2024c) employs a bidirectional Mamba encoder block to capture inter-variate correlations and a feed forward network temporal dependency encoding layer to learn temporal sequence dynamics.
 This novel approach is the current SOTA model for TSF task and forms the foundation of our proposed method.
-  • iTransformer [26] inverts the order of sequence processing by first analyzing each individual variate separately and then merging the information across all variates.
-  • PatchTST [28] divides the time-series into sub-series patches treated as input tokens. It leverages channel-independent shared embeddings and weights for efficient representation learning.
-  • Crossformer [46] embeds multivariate time-series into a 2D vector array preserving time and dimension information and introduces two-stage attention to capture both cross-time and cross-dimension dependencies.
-  • FEDformer [49] is a frequency-enhanced Transformer that uses trend-seasonality decomposition and exploit sparse representations, Fourier transform, of time-series data to achieve linear complexity to sequence length.
-  • Autoformer [39] constructs a decomposition architecture that employs traditional sequence decomposition in its inner blocks and utilizes an auto-correlation mechanism.
-  • DLinear [44] maps trend and seasonality components into predictions via a single linear layer.
-  • TiDE [9] is a MLP based encoder-decoder model that is best suitable for linear dynamical systems.
-  • RLinear [23] is the current SOTA linear model that introduces reversible normalization and channel independence within a purely linear structure.
-  • TimesNet [40] proposes a task-general backbone, TimesBlock, that transforms 1D time-series into 2D tensors and uses 2D convolution kernels to capture intra-period and inter-period variations.
+  • 2.
+    iTransformer Liu et al. (2023) inverts the order of sequence processing by first analyzing each individual variate separately and then merging the information across all variates.
+  • 3.
+    PatchTST Nie et al.
+(2022) divides the time-series into sub-series patches treated as input tokens.
+It leverages channel-independent shared embeddings and weights for efficient representation learning.
+  • 4.
+    Crossformer Zhang and Yan (2023) embeds multivariate time-series into a 2D vector array preserving time and dimension information and introduces two-stage attention to capture both cross-time and cross-dimension dependencies.
+  • 5.
+    FEDformer Zhou et al.
+(2022) is a frequency-enhanced Transformer that uses trend-seasonality decomposition and exploit sparse representations, Fourier transform, of time-series data to achieve linear complexity to sequence length.
+  • 6.
+    Autoformer Wu et al. (2021) constructs a decomposition architecture that employs traditional sequence decomposition in its inner blocks and utilizes an auto-correlation mechanism.
+  • 7.
+    DLinear Zeng et al. (2023) maps trend and seasonality components into predictions via a single linear layer.
+  • 8.
+    TiDE Das et al. (2023) is a MLP based encoder-decoder model that is best suitable for linear dynamical systems.
+  • 9.
+    RLinear Li et al. (2023) is the current SOTA linear model that introduces reversible normalization and channel independence within a purely linear structure.
+  • 10.
+    TimesNet Wu et al. proposes a task-general backbone, TimesBlock, that transforms 1D time-series into 2D tensors and uses 2D convolution kernels to capture intra-period and inter-period variations.
 Training and Implementation Details. See the Suppl Mat for training and implementation details, especially the hyperparameters.
-Performance Measure. Following the common practice (e.g., [34, 26, 28]), models’ performances are compared using the Mean Square Error (MSE) as defined in Equation (12).
-Models ms-Mamba ms-Mamba w/ Learnable $\Delta$ S-Mamba iTrans- former RLinear PatchTST Cross- former TiDE TimesNet DLinear FED- former Auto- former
-ETTm1 96 0.328 0.326 0.333 0.334 0.355 0.329 0.404 0.364 0.338 0.345 0.379 0.505
-192 0.372 0.372 0.376 0.377 0.391 0.367 0.450 0.398 0.374 0.380 0.426 0.553
-336 0.406 0.406 0.408 0.426 0.424 0.399 0.532 0.428 0.410 0.413 0.445 0.621
-720 0.470 0.470 0.475 0.491 0.487 0.454 0.666 0.487 0.478 0.474 0.543 0.671
-Avg 0.394 0.394 0.398 0.407 0.414 0.387 0.513 0.419 0.400 0.403 0.448 0.588
-ETTm2 96 0.176 0.175 0.179 0.180 0.182 0.175 0.287 0.207 0.187 0.193 0.203 0.255
-192 0.244 0.244 0.250 0.250 0.246 0.241 0.414 0.290 0.249 0.284 0.269 0.281
-336 0.309 0.306 0.312 0.311 0.307 0.305 0.597 0.377 0.321 0.369 0.325 0.339
-720 0.408 0.407 0.411 0.412 0.407 0.402 1.730 0.558 0.408 0.554 0.421 0.433
-Avg 0.284 0.283 0.288 0.288 0.286 0.281 0.757 0.358 0.291 0.350 0.305 0.327
-ETTh1 96 0.384 0.384 0.386 0.386 0.386 0.414 0.423 0.479 0.384 0.386 0.376 0.449
-192 0.437 0.438 0.443 0.441 0.437 0.460 0.471 0.525 0.435 0.436 0.420 0.500
-336 0.479 0.482 0.489 0.487 0.479 0.501 0.570 0.565 0.491 0.481 0.459 0.521
-720 0.482 0.493 0.502 0.503 0.481 0.500 0.653 0.594 0.521 0.519 0.506 0.514
-Avg 0.445 0.449 0.455 0.454 0.446 0.469 0.529 0.541 0.458 0.456 0.440 0.496
-ETTh2 96 0.291 0.291 0.296 0.297 0.288 0.302 0.745 0.400 0.340 0.333 0.358 0.346
-192 0.371 0.369 0.376 0.380 0.374 0.388 0.877 0.528 0.402 0.477 0.429 0.456
-336 0.411 0.412 0.424 0.428 0.415 0.426 1.043 0.643 0.452 0.594 0.496 0.482
-720 0.418 0.418 0.426 0.427 0.420 0.431 1.104 0.874 0.462 0.831 0.463 0.515
-Avg 0.373 0.373 0.381 0.383 0.374 0.387 0.942 0.611 0.414 0.559 0.437 0.450
-Table 2: Experiment 1: Quantitative comparison between ms-Mamba and the existing methods on ETT Datasets.
+Performance Measure.
+Following the common practice (e.g., Wang et al.
+(2024c); Liu et al.
+(2023); Nie et al.
+(2022)), models’ performances are compared using the Mean Square Error (MSE) as defined in Equation (12).
+Table 2: Experiment 1: Quantitative results on ETT Datasets.
 The lookback length $L$ is set to 96 and the forecast length $T$ is set to 96, 192, 336, 720.
 Top results are highlighted in bold while the second bests are underlined.
-Models ms-Mamba ms-Mamba w/ Learnable $\Delta$ S-Mamba iTrans- former RLinear PatchTST Cross- former TiDE TimesNet DLinear FED- former Auto- former
-Electricity 96 0.137 0.138 0.139 0.148 0.201 0.181 0.219 0.237 0.168 0.197 0.193 0.201
-192 0.157 0.157 0.159 0.162 0.201 0.188 0.231 0.236 0.184 0.196 0.201 0.222
-336 0.171 0.174 0.176 0.178 0.215 0.204 0.246 0.249 0.198 0.209 0.214 0.231
-720 0.195 0.199 0.204 0.225 0.257 0.246 0.280 0.284 0.220 0.245 0.246 0.254
-Avg 0.165 0.167 0.170 0.178 0.219 0.205 0.244 0.251 0.192 0.212 0.214 0.227
-Exchange 96 0.085 0.086 0.086 0.086 0.093 0.088 0.256 0.094 0.107 0.088 0.148 0.197
-192 0.177 0.178 0.182 0.177 0.184 0.176 0.470 0.184 0.226 0.176 0.271 0.300
-336 0.325 0.326 0.332 0.331 0.351 0.301 1.268 0.349 0.367 0.313 0.460 0.509
-720 0.843 0.843 0.867 0.847 0.886 0.901 1.767 0.852 0.964 0.839 1.195 1.447
-Avg 0.358 0.358 0.367 0.360 0.378 0.367 0.940 0.370 0.416 0.354 0.519 0.613
-Weather 96 0.163 0.163 0.165 0.174 0.192 0.177 0.158 0.202 0.172 0.196 0.217 0.266
-192 0.213 0.213 0.214 0.221 0.240 0.225 0.206 0.242 0.219 0.237 0.276 0.307
-336 0.269 0.270 0.274 0.278 0.292 0.278 0.272 0.287 0.280 0.283 0.339 0.359
-720 0.349 0.349 0.350 0.358 0.364 0.354 0.398 0.351 0.365 0.345 0.403 0.419
-Avg 0.249 0.249 0.251 0.258 0.272 0.259 0.259 0.271 0.259 0.265 0.309 0.338
-Solar Ener. 96 0.196 0.195 0.205 0.203 0.322 0.234 0.310 0.312 0.250 0.290 0.242 0.884
-192 0.230 0.230 0.237 0.233 0.359 0.267 0.734 0.339 0.296 0.320 0.285 0.834
-336 0.250 0.247 0.258 0.248 0.397 0.290 0.750 0.368 0.319 0.353 0.282 0.941
-720 0.249 0.249 0.260 0.249 0.397 0.289 0.769 0.370 0.338 0.356 0.357 0.882
-Avg 0.231 0.229 0.240 0.233 0.369 0.270 0.641 0.347 0.301 0.330 0.291 0.885
-Table 3: Experiment 1: Quantitative comparison between ms-Mamba and the existing methods on Electricity, Exchange, Weather and Solar-Energy Datasets.
+Dataset $T$ ms-Mamba (Ours) S-Mamba iTransformer RLinear PatchTST Crossformer TiDE TimesNet DLinear FEDformer Autoformer
+ETTm1 96 0.326 0.333 0.334 0.355 0.329 0.404 0.364 0.338 0.345 0.379 0.505
+192 0.371 0.376 0.377 0.391 0.367 0.450 0.398 0.374 0.380 0.426 0.553
+336 0.406 0.408 0.426 0.424 0.399 0.532 0.428 0.410 0.413 0.445 0.621
+720 0.470 0.475 0.491 0.487 0.454 0.666 0.487 0.478 0.474 0.543 0.671
+[.4pt/1pt] Avg 0.394 0.398 0.407 0.414 0.387 0.513 0.419 0.400 0.403 0.448 0.588
+ETTm2 96 0.175 0.179 0.180 0.182 0.175 0.287 0.207 0.187 0.193 0.203 0.255
+192 0.244 0.250 0.250 0.246 0.241 0.414 0.290 0.249 0.284 0.269 0.281
+336 0.306 0.312 0.311 0.307 0.305 0.597 0.377 0.321 0.369 0.325 0.339
+720 0.407 0.411 0.412 0.407 0.402 1.730 0.558 0.408 0.554 0.421 0.433
+[.4pt/1pt] Avg 0.283 0.288 0.288 0.286 0.281 0.757 0.358 0.291 0.350 0.305 0.327
+ETTh1 96 0.384 0.386 0.386 0.386 0.414 0.423 0.479 0.384 0.386 0.376 0.449
+192 0.438 0.443 0.441 0.437 0.460 0.471 0.525 0.435 0.436 0.420 0.500
+336 0.482 0.489 0.487 0.479 0.501 0.570 0.565 0.491 0.481 0.459 0.521
+720 0.493 0.502 0.503 0.481 0.500 0.653 0.594 0.521 0.519 0.506 0.514
+[.4pt/1pt] Avg 0.449 0.455 0.454 0.446 0.469 0.529 0.541 0.458 0.456 0.440 0.496
+ETTh2 96 0.291 0.296 0.297 0.288 0.302 0.745 0.400 0.340 0.333 0.358 0.346
+192 0.369 0.376 0.380 0.374 0.388 0.877 0.528 0.402 0.477 0.429 0.456
+336 0.412 0.424 0.428 0.415 0.426 1.043 0.643 0.452 0.594 0.496 0.482
+720 0.418 0.426 0.427 0.420 0.431 1.104 0.874 0.462 0.831 0.463 0.515
+[.4pt/1pt] Avg 0.373 0.381 0.383 0.374 0.387 0.942 0.611 0.414 0.559 0.437 0.450
+Table 3: Experiment 1: Quantitative results on Electricity, Exchange, Weather and Solar-Energy Datasets.
 The lookback length $L$ is set to 96 and the forecast length $T$ is set to 96, 192, 336, 720.
 Top results are highlighted in bold while the second bests are underlined.
+Dataset $T$ ms-Mamba (Ours) S-Mamba iTransformer RLinear PatchTST Crossformer TiDE TimesNet DLinear FEDformer Autoformer
+Electricity 96 0.138 0.139 0.148 0.201 0.181 0.219 0.237 0.168 0.197 0.193 0.201
+192 0.157 0.159 0.162 0.201 0.188 0.231 0.236 0.184 0.196 0.201 0.222
+336 0.174 0.176 0.178 0.215 0.204 0.246 0.249 0.198 0.209 0.214 0.231
+720 0.199 0.204 0.225 0.257 0.246 0.280 0.284 0.220 0.245 0.246 0.254
+[.4pt/1pt] Avg 0.167 0.170 0.178 0.219 0.205 0.244 0.251 0.192 0.212 0.214 0.227
+Exchange 96 0.086 0.086 0.086 0.093 0.088 0.256 0.094 0.107 0.088 0.148 0.197
+192 0.178 0.182 0.177 0.184 0.176 0.470 0.184 0.226 0.176 0.271 0.300
+336 0.326 0.332 0.331 0.351 0.301 1.268 0.349 0.367 0.313 0.460 0.509
+720 0.843 0.867 0.847 0.886 0.901 1.767 0.852 0.964 0.839 1.195 1.447
+[.4pt/1pt] Avg 0.358 0.367 0.360 0.378 0.367 0.940 0.370 0.416 0.354 0.519 0.613
+Weather 96 0.163 0.165 0.174 0.192 0.177 0.158 0.202 0.172 0.196 0.217 0.266
+192 0.213 0.214 0.221 0.240 0.225 0.206 0.242 0.219 0.237 0.276 0.307
+336 0.270 0.274 0.278 0.292 0.278 0.272 0.287 0.280 0.283 0.339 0.359
+720 0.349 0.350 0.358 0.364 0.354 0.398 0.351 0.365 0.345 0.403 0.419
+[.4pt/1pt] Avg 0.249 0.251 0.258 0.272 0.259 0.259 0.271 0.259 0.265 0.309 0.338
+Sol. Ener. 96 0.195 0.205 0.203 0.322 0.234 0.310 0.312 0.250 0.290 0.242 0.884
+192 0.230 0.237 0.233 0.359 0.267 0.734 0.339 0.296 0.320 0.285 0.834
+336 0.247 0.258 0.248 0.397 0.290 0.750 0.368 0.319 0.353 0.282 0.941
+720 0.249 0.260 0.249 0.397 0.289 0.769 0.370 0.338 0.356 0.357 0.882
+[.4pt/1pt] Avg 0.229 0.240 0.233 0.369 0.270 0.641 0.347 0.301 0.330 0.291 0.885
 5.2 Experiment 1: Comparison with State-of-the-Art
-We compare ms-Mamba with fixed and learnable temporal scales against SOTA methods over 13 benchmark datasets.
+Unless stated otherwise, ms-Mamba refers to ms-Mamba with learnable temporal scales.
 Traffic-related Datasets (Table 1).
-The results in Table 1 show that ms-Mamba with fixed or learnable temporal scales provides the best or second best results over all traffic datasets across all forecast lengths.
-Compared to our baseline model of S-Mamba, ms-Mamba delivers significant improvements, especially on the Traffic dataset.
+The results in Table 1 show that ms-Mamba provides the best or second best results over all traffic datasets across all forecast lengths.
+Compared to our baseline model of S-Mamba, ms-Mamba delivers significant improvements, especially on the Traffic dataset (0.402 vs.
+0.414).
 ETT Datasets (Table 2).
-On ETT datasets, as in Table 2, ms-Mamba with fixed or learnable temporal scales is typically the second best and the best performing method in ETTh2 and in some configurations of ETTm1 and ETTm2.
+On ETT datasets, as in Table 2, ms-Mamba is typically the second best and the best performing method in ETTh2 and in some configurations of ETTm1 and ETTm2.
 Compared to S-Mamba, ms-Mamba provides better performance.
 Other Datasets (Table 3).
-The results in other datasets (Table 3) are in agreement with the results on Traffic-related and ETT datasets: our ms-Mamba with fixed or learnable temporal scales performs better than or is generally on par with SOTA methods, and consistently provides better performance than the S-Mamba baseline.
-                                    Dataset $\Rightarrow$ Traffic ETTh1 Solar Energy
-                             VC Encoding $\Rightarrow$ 96 192 336 720 96 192 336 720 96 192 336 720
-ms-Mamba with fixed scales: $\mathbf{\alpha}=(1,2,4,8)$ 0.376 0.392 0.405 0.452 0.384 0.437 0.479 0.482 0.196 0.230 0.250 0.249
-ms-Mamba with fixed scales: $\mathbf{\alpha}=(0.5,1,1.5,2)$ 0.374 0.389 0.414 0.458 0.386 0.438 0.481 0.491 0.197 0.232 0.248 0.251
-ms-Mamba with fixed scales: $\mathbf{\alpha}=(1,2,3,4)$ 0.390 0.403 0.415 0.455 0.386 0.439 0.482 0.495 0.196 0.232 0.250 0.251
-ms-Mamba with fixed scales: $\mathbf{\alpha}=(1,4,8,16)$ 0.380 0.411 0.421 0.453 0.385 0.439 0.484 0.495 0.197 0.232 0.250 0.251
-ms-Mamba with learnable scales 0.375 0.384 0.408 0.442 0.384 0.438 0.482 0.493 0.195 0.230 0.247 0.249
-ms-Mamba with dynamic scales 0.376 0.390 0.414 0.454 0.384 0.440 0.480 0.493 0.194 0.231 0.249 0.251
-Table 4: Experiment 2: Ablation study on Traffic, ETTh1 and Solar Energy datasets (one dataset from each category).
+The results in other datasets (Table 3) are in agreement with the results on Traffic-related and ETT datasets: Our ms-Mamba performs better than or is generally on par with SOTA methods, and consistently provides better performance than the S-Mamba baseline, especially on the Solar Energy dataset (0.229 vs.
+0.240).
+Summary: Our ms-Mamba exhibits strong performance across 13 benchmark datasets and diverse problems compared to the baseline method of s-Mamba and the SOTA methods.
+We observe that the performance gains are particularly pronounced on datasets with distinct, hierarchical temporal patterns (e.g., Traffic and Solar Energy).
+This supports our hypothesis that assigning distinct Mamba blocks to different sampling rates allows the model to better disentangle and capture these overlapping temporal dynamics compared to single-scale baseline.
+Table 4: Experiment 2: Ablation study on Traffic and Solar Energy datasets.
 The lookback length $L=96$, while the forecast length $T\in\left\{96,192,336,720\right\}$.
-$\mathbf{\alpha}=(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4})$ indicates that the $\Delta_{1}$ (learnable sampling rate of the base Mamba) is multiplied by these coefficients to obtain the sampling rates for the Mamba blocks.
+$\mathbf{\alpha}=(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4})$ indicates that the $\Delta_{1}$ (learnable sampling rate of the base Mamba) is multiplied by these coefficients to obtain the sampling rates for the Mamba blocks (Eq.
+8).
 Top results are highlighted in bold while the second bests are underlined.
+     Dataset $\Rightarrow$ Traffic Solar Energy
+Prediction Length $\Rightarrow$ 96 192 336 720 96 192 336 720
+ms-Mamba with fixed temporal scales
+$\mathbf{\alpha}=(1,2,4,8)$ .376 .392 .405 .452 .196 .230 .250 .249
+$\mathbf{\alpha}=(0.5,1,1.5,2)$ .374 .389 .414 .458 .197 .232 .248 .251
+$\mathbf{\alpha}=(1,2,3,4)$ .390 .403 .415 .455 .196 .232 .250 .251
+$\mathbf{\alpha}=(1,4,8,16)$ .380 .411 .421 .453 .197 .232 .250 .251
+ms-Mamba with learnable scales
+.375 .384 .408 .442 .195 .230 .247 .249
+ms-Mamba with dynamic scales
+.376 .390 .414 .454 .194 .231 .249 .251
+Table 5: Experiment 2: Effect of the number of scales using the Solar Energy dataset with $T=96$.
+Method / Scale Count 2 3 4 5 6
+ms-Mamba 0.202 0.199 0.196 0.199 0.203
 5.3 Experiment 2: Ablation Analysis
-In this experiment, we investigate the different strategies for integrating multiple-scales into ms-Mamba.
-For this, we evaluate the performances of the strategies described in Section 4.2: (i) ms-Mamba with fixed temporal scales, where we multiply the learnable sampling rate $\Delta_{1}$ by fixed hyperparameters, $\Delta_{i}=\alpha_{i}\times\Delta_{1}$.
-(ii) ms-Mamba with learnable temporal scales, where each $\Delta_{i}$ is a learnable parameter.
-(iii) ms-Mamba with dynamic temporal scales, where each $\Delta_{i}$ is estimated by an MLP applied on the input embeddings.
+Evaluation of Different Strategies for Setting $\Delta_{i}$.
+First, we evaluate the performances of the strategies described in Section 4.2: (i) ms-Mamba with fixed temporal scales.
+(ii) ms-Mamba with learnable temporal scales.
+(iii) ms-Mamba with dynamic temporal scales.
 For this analysis, we tune the hyperparameters in all settings.
 To keep the number of experiments manageable, we consider one dataset from each category.
-The results in Table 4 suggest that using fixed temporal scales with $\alpha_{i}$ coefficients of $(1,2,4,8)$ performs best among different coefficients considered (6 best and 2 second-best results out of 12 datasets).
-Smaller or larger coefficients than $(1,2,4,8)$ do not appear to provide better results overall.
-Moreover, ms-Mamba with learnable $\Delta_{i}$ provides slightly better results than the best fixed temporal scale version (provides 6 best and 4 second-best results).
-Fixed temporal scales introduces several hyperparameters to be tuned, which is a significant limitation.
-Therefore, ms-Mamba with learnable temporal scales can be the preferred option.
-However, in Experiment 1, we have considered both versions for a more extensive evaluation.
-Models ms-Mamba S-Mamba
+The results in Table 4 suggest that using learned temporal scales performs best among the different strategies considered (5 best and 3 second-best results out of 8 settings).
+Fixed temporal scales provide on par results in many different settings.
+These results suggest that a more refined search for fixed temporal scales might provide better results.
+However, this requires more experiments for hyperparameter tuning, which is alleviated by the learnable temporal scales approach.
+The dynamic scales strategy generally provides the worst performance, except for the Solar Energy dataset.
+The subpar performance of the dynamic approach may be owing to the extra learnable parameters introduced by the approach.
+Effect of the number of scales.
+In this analysis, for the best performing ms-Mamba version (ms-Mamba with temporal scales), we evaluate the impact of the number of scales considered.
+As listed in Table 5, we observe that having 4 scales provides the best performance overall.
+Effect of input length.
+In Figure 3, we analyze the impact of the input length on the performance of ms-Mamba.
+We see that larger input length provides better results for the Solar and Electricity datasets whereas an input length of 96 provides the best results for the ETTh2 dataset.
+Although there is no single input length that works best for all datasets, for fair comparison with the literature, we have used 96 as the input length in all experiments.
+Figure 3: Effect of input length ($L$) on the performance of ms-Mamba.
+Summary: Learnable temporal scales provides the best results whilst not requiring any hyperparameter tuning as in the fixed temporal scales approach.
+Therefore, unless stated otherwise, ms-Mamba refers to ms-Mamba with learnable temporal scales.
+Figure 4: Experiment 3: Values of learnable $\Delta_{i}$ parameters over time for the Solar Energy dataset where $T=96$ ($\Delta_{i}$ parameters are initialized randomly between 1 and 4).
+(a) ms-Mamba
+(b) S-Mamba
+Figure 5: Experiment 3: Forecast comparison between ms-Mamba and S-Mamba on four datasets when the input length is 96 ($L=96$) and the forecast length is 96 ($T=96$).
+The green line represents the input, the blue line represents the ground truth and the red line represents the forecast.
+Table 6: Experiment 4: Efficiency comparison of ms-Mamba and S-Mamba.
+The lookback length $L$ is set to 96 and the forecast length $T$ is set to 96, 192, 336, 720.
+We report results on the extended ETT dataset family to demonstrate intra-category consistency, alongside representative datasets from other categories.
+Models ms-Mamba (Ours) S-Mamba
 Metric MSE #Params Memory MACs MSE #Params Memory MACs
+ETTh1 96 0.384 0.195M 0.746MB 0.063G 0.386 1.145M 4.37MB 0.400G
+192 0.438 0.179M 0.685MB 0.057G 0.443 1.170M 4.46MB 0.409G
+336 0.482 0.270M 1.03MB 0.087G 0.489 1.207M 4.60MB 0.422G
+720 0.493 0.223M 0.85MB 0.075G 0.502 1.306M 4.98MB 0.457G
 ETTh2 96 0.291 0.481M 1.84MB 0.165G 0.296 1.150M 4.40MB 1.563G
@@
 720 0.418 0.552M 2.11MB 0.195G 0.426 1.311M 5.00MB 1.675G
-Solar Ener. 96 0.195 3.958M 15.10MB 16.72G 0.205 4.643M 17.71MB 20.00G
+ETTm1 96 0.326 0.160M 0.61MB 0.055G 0.333 1.15M 4.37MB 0.400G
+192 0.371 0.121M 0.46MB 0.028G 0.376 1.201M 17.90MB 0.109G
+336 0.406 0.191M 0.73MB 0.065G 0.408 0.333M 1.272B 0.116G
+720 0.470 0.223M 0.85MB 0.075G 0.475 0.383M 1.46MB 0.133G
+ETTm2 96 0.175 0.304M 1.16MB 0.102G 0.179 1.15M 4.37MB 0.400G
+192 0.244 0.168M 0.64MB 0.058G 0.250 0.315M 1.20MB 0.109G
+336 0.306 0.132M 0.505MB 0.044G 0.312 0.333M 1.272B 0.116G
+720 0.407 0.189M 0.72MB 0.063G 0.411 0.383M 1.46MB 0.133G
+Solar En. 96 0.195 3.958M 15.10MB 16.72G 0.205 4.643M 17.71MB 20.00G
 192 0.230 2.028M 7.74MB 8.57G 0.237 4.692M 17.90MB 20.21G
@@
 720 0.442 29.68M 114.5MB 407.9G 0.460 9.507M 36.26MB 129.5G
-Table 5: Experiment 3: Performance comparison of ms-Mamba and S-Mamba on one dataset from each category. The lookback length $L$ is set to 96 and the forecast length $T$ is set to 96, 192, 336, 720.
-5.4 Experiment 3: Efficiency Analysis
-In this experiment, we investigate the efficiency of ms-Mamba (with learnable temporal scales, as it provides the best results and incurs more learnable parameters) in comparison with the baseline S-Mamba.
-In each dataset and for each method, we provide the results of the best performing configurations.
-As listed in Table 5, on the ETTh1 and Solar Energy datasets, we see that ms-Mamba provides the best results with less parameters, memory and operations.
+5.4 Experiment 3: Qualitative Analysis
+Evolution of learned $\alpha$.
+In Figure 4, we investigate how the learned $\Delta_{i}$ changes throughout training.
+We observe that the $\Delta_{i}$ values fluctuate during the initial stages of training.
+However, they converge to and saturate at certain values.
+[…]
+5.5 Experiment 4: Efficiency Analysis
+In this experiment, we investigate the efficiency of ms-Mamba (with learnable temporal scales) in comparison with the baseline s-Mamba.
+In each dataset, we provide the results of the best performing configurations for each method.
+As listed in Table 6, on the datasets from ETT category and Solar Energy dataset, we see that ms-Mamba provides the best results with less parameters, memory and operations.
 This is crucial as it shows multiple temporal scales can be utilized with less computational overhead.
 However, this result is not observed in the Traffic dataset because it contains significantly more variates (862) compared to the ETTh2 (7) and Solar Energy (137) datasets.
-Although ms-Mamba provides better results with more variates, it is not able to do so with less parameters, memory and computations as in ETTh2 and Solar Energy datasets.
+Summary: ms-Mamba provides better performance than the baseline S-Mamba with less parameters, memory and computations on many datasets. However, this is not observed for the Traffic dataset.
 6 Conclusion
@@
 Our architecture extends Mamba (or its derivative S-Mamba) where we include several Mamba blocks with different sampling rates to process multiple temporal scales simultaneously.
-The different sampling rates can either be fixed or learned from the data, which leads to a simple architecture with a multi-scale ability.
-Our results on several TSF benchmarks show that our approach provides the best or on par performance compared to SOTA methods.
-What is remarkable is that, compared to the baseline model (S-Mamba), ms-Mamba provides better results with less parameters, memory, and operations.
+The different sampling rates can either be fixed or learned from the data, which leads to a simple architecture with a multi-scale processing ability.
+Our results on 13 TSF benchmarks over different problem kinds show that our approach provides the best or on par performance compared to the state-of-the-art methods.
+What is remarkable is that, compared to the baseline model (S-Mamba), ms-Mamba provides better results and, on many datasets, does so with less parameters, memory, and operations.
 Limitations and Future Work.
@@
 Moreover, ms-Mamba can complement other types of deep modules, e.g., scaled-dot-product attention, linear attention.
+Furthermore, while our averaging fusion was effective and efficient, exploring more sophisticated, learnable fusion mechanisms (e.g., attention or weighted-averaging) to better exploit the complementarity between scales is a promising direction for future work.
+Acknowledgment
+This work has been partially funded by the European Union’s Horizon project LEGOFIT - Adaptable technological solutions based on early design actions for the construction and renovation of Energy Positive Homes (Grant agreement ID: 101104058).
+We also gratefully acknowledge the computational resources kindly provided by METU Image Lab and METU-ROMER, Center for Robotics and Artificial Intelligence.
+[…]
+By employing multiple Mamba blocks, each parameterized with a different sampling rate $\Delta_{i}$, our model is designed to intrinsically capture features at multiple, distinct temporal resolutions (or time scales).
+Recall from Section 3.2 that:
+$\hat{A}$ $=\exp(\Delta A),$ (13)
+$\hat{B}$ $=A^{-1}(\exp(\Delta A)-I)\,B\approx(\Delta A)^{-1}(\exp(\Delta A)-I)\,\Delta B,$ (14)
+Thus, the choice of $\Delta$ directly controls the discrete-time dynamics by altering both the transition matrix $\hat{A}$ and the input-response matrix $\hat{B}$.
+Using multiple SSMs with different $\Delta$ therefore induces multiple characteristic timescales in the latent dynamics.
+A.1 Influence of Small $\Delta$
+When $\Delta$ is small, the discretization stays close to the identity map:
+$\hat{A}=\exp(\Delta A)\approx I+\Delta A.$ (15)
+This implies that the eigenvalues of $\hat{A}$ lie close to 1, leading to slow decay of the hidden state.
+Consequently, the SSM retains information over many discrete time steps, capturing fine-grained temporal variations and high-frequency structure.
+The input mapping behaves similarly:
+$\hat{B}\approx\Delta B,$ (16)
+so the effective input contribution per time step is small but frequent.
+Taken together, small $\Delta$ yields a high-resolution temporal process: the latent state evolves slowly, and the model is sensitive to rapid changes in the input.
+A.2 Influence of Large $\Delta$
+For large $\Delta$, the discretized transition becomes
+$\hat{A}=\exp(\Delta A),$ (17)
+which strongly contracts the dynamics whenever $A$ has negative real eigenvalues.
+Eigenvalues of $\exp(\Delta A)$ shrink rapidly in magnitude as $\Delta$ increases, causing fast decay of the hidden state.
+This yields a coarse-grained temporal process: the model integrates information over broad temporal windows and captures slow trends rather than rapid fluctuations.
+The behavior of the input operator reflects this as well. Since in the following definition:
+$\hat{B}=A^{-1}(\exp(\Delta A)-I)B,$ (18)
+large $\Delta$ causes $(\exp(\Delta A)-I)$ to amplify modes of $A$ with large negative real parts, producing a stronger input influence per update step.
+This leads the model to respond primarily to aggregated (low-frequency) input structure.
+Overall, large $\Delta$ induces shorter memory and lower temporal resolution, complementing the long-memory, high-resolution behavior produced by small $\Delta$.
+Appendix B More Details about the Experimental Setup
+B.1 Dataset Statistics
+In this section (in Table 7), we summarize the properties of the datasets used in the experiments.
+Table 7: 13 public benchmark datasets used in the paper.
+Datasets Variates Timesteps Resolution
+Traffic 862 17,544 1 hour
+PEMS03 358 26,209 5 minutes
+PEMS04 307 16,992 5 minutes
+PEMS07 883 28,224 5 minutes
+PEMS08 170 17,856 5 minutes
+ETTm1 & ETTm2 7 17,420 15 minutes
+ETTh1 & ETTh2 7 69,680 1 hour
+Electricity 321 26,304 1 hour
+Exchange 8 7,588 1 day
+Weather 21 52,696 10 minutes
+Solar Energy 137 52,560 10 minutes
+B.2 Training and Implementation Details
+For a fair comparison, we follow the experimental settings of S-Mamba Wang et al.
+(2024c) where input sequence length is fixed at 96 time steps and prediction lengths for training and testing are fixed at 96, 192, 336 and 720 for each dataset.
+To ensure a fair comparison, we adopt the specific epoch settings from the S-Mamba baseline: the maximum number of training epochs is set to 5 for PEMS03, Electricity, and Weather, and 10 for the remaining datasets (Traffic, PEMS04, PEMS07, PEMS08, the ETT category, Exchange, and Solar Energy).
+We tuned the learning rate, batch size, the number of encoder layers, embedding dimension and FFN dimension in our experiments – see Table 8 for the ranges considered.
+For datasets with fewer variates, we explored smaller embedding and FFN dimensions, whereas for datasets with more variates, we explored higher dimensions.
+For alignment of results with the original experimental set-up, we did not explicitly tune the random seed.
+During the hyperparameter optimization phase, we measure the mean-squared error (MSE) performance on the validation set and use early stopping.
+ms-Mamba models are implemented in PyTorch framework.
+Our experiments are conducted on a cluster containing NVIDIA RTX 3090 and NVIDIA RTX A6000 GPUs.
+Table 8: Model Hyperparameters and Their Ranges
+Hyperparameter Range
+Learning Rate ($1e^{-5}$, $1e^{-3}$)
+Batch Size {16, 32, 64}
+Encoder Layers {1, 2, 4}
+Embedding Dimension {32, 64, 80, 128, 256, 512, 768, 1024}
+FFN Dimension {32, 64, 80, 128, 256, 512, 768, 1024}
+B.3 Additional Visualizations of Multi-scale Patterns
+To further substantiate the motivation presented in the Introduction, we provide additional visualizations from different domains for one dataset in the ETT category (Figure 6(a)) and the Traffic dataset (Figure 6(b)).
+These plots demonstrate that the multi-scale nature of time-series data where signal characteristics change across different temporal resolutions is a ubiquitous phenomenon, not limited to weather data.
+(a) ETTh1 dataset
+(b) Traffic dataset
+Figure 6: Multi-scale visualization for (a) the ETTh1 dataset from ETT category and (b) Traffic dataset.
+The original signal (red) contains high-frequency fluctuations, while the smoothed versions reveal underlying and longer-term trends, illustrating the necessity of capturing dynamics at multiple resolutions.